### v1

# On Hardness and Approximation of Broadcasting in Sparse Graphs

Abstract
We study the Telephone Broadcasting problem in sparse graphs. Given a designated source in an undirected graph, the task is to disseminate a message to all vertices in the minimum number of rounds, where in each round every informed vertex may inform at most one uninformed neighbor. For general graphs with $n$ vertices, the problem is NP-hard. Recent work shows that the problem remains NP-hard even on restricted graph classes such as cactus graphs of pathwidth $2$ [Aminian et al., ICALP 2025] and graphs at distance-1 to a path forest [Egami et al., MFCS 2025].
In this work, we investigate the problem in several sparse graph families. We first prove NP-hardness for $k$-cycle graphs, namely graphs formed by $k$ cycles sharing a single vertex, as well as $k$-path graphs, namely graphs formed by $k$ paths with shared endpoints. Despite multiple efforts to understand the problem in these simple graph families, the computational complexity of the problem had remained unsettled, and our hardness results answer open questions by Bhabak and Harutyunyan [CALDAM 2015] and Harutyunyan and Hovhannisyan [COCAO 2023] concerning the problem’s complexity in $k$-cycle and $k$-path graphs, respectively.
On the positive side, we present Polynomial-Time Approximation Schemes (PTASs) for $k$-cycle and $k$-path graphs, improving over the best existing approximation factors of $2$ for $k$-cycle graphs and an approximation factor of $4$ for $k$-path graphs. Moreover, we identify a structural frontier for tractability by showing that the problem is solvable in polynomial time on graphs of bounded bandwidth. This result generalizes existing tractability results for special sparse families such as necklace graphs.
1 Introduction
In the Telephone Broadcasting problem, the objective is to spread a message from a designated source vertex to all vertices of a network via a sequence of telephone calls. The network is modeled as an undirected, unweighted graph on $n$ vertices, and communication proceeds in synchronous rounds. Initially, only the source knows the message. In each round, every informed vertex may contact at most one uninformed neighbor to pass along the message. The goal is to minimize the total number of rounds until all vertices are informed.
The problem is known to be computationally difficult. Slater et al.  [19] proved that Telephone Broadcasting is NP-complete. Elkin and Kortsarz [7] established that approximating Telephone Broadcasting within a factor of $3-\epsilon$ for any $\epsilon>0$ is NP-hard. On the other hand, Elkin and Kortsarz [8] designed an $\mathcal{O}(\log n/\log\log n)$-approximation algorithm for general graphs, which is the best existing approximation algorithm. Whether a constant-factor approximation exists for general graphs remains open, though such results are known for specific classes, including unit disk graphs [18] and cactus graphs [1].
A substantial body of work has investigated the complexity of Telephone Broadcasting on restricted graph classes. On the positive side, polynomial-time algorithms are known for several families, including trees [10], grids, stars of cliques, fully connected trees, chains of rings, necklace graphs, and certain subclasses of cactus graphs; see, e.g., [5, 12, 17, 20, 9]. On the negative side, hardness results persist even for structurally simple graphs: the problem is NP-hard in planar cubic graphs and in graphs of bounded treewidth. More recently, Tale showed NP-hardness for graphs of pathwidth $3$, while Aminian et al. [1] established hardness for so-called snowflake graphs (cactus graphs of pathwidth $2$), resolving a long-standing open question about the complexity of broadcasting in cacti. Even stronger results were obtained by Egami et al. [6], who proved NP-hardness for graphs at distance 1 from path forests, i.e., graphs in which deleting a single vertex yields a path forest. See Figure 1 for illustrations of snowflake graphs and graphs at distance 1 from path forests.
These results highlight intriguing frontier cases where the complexity of the problem remains unresolved. One such case is that of $k$-cycle graphs (also called flower graphs), obtained by joining $k$ cycles at a single vertex. Despite their simple definition, these graphs exhibit surprisingly rich behavior. Broadcasting on $k$-cycle graphs has been studied from an approximation perspective: the best known algorithm achieves a factor of 2 [2], yet the exact complexity remains open. A closely related family is that of $k$-path graphs, formed by joining $k$ paths at their endpoints. Several approximation algorithms are known for $k$-path graphs [3, 14], with the current best result being a 2-approximation [14]. Nevertheless, the precise computational complexity of the problem for this family also remains open.
1.1 Contribution
Our contributions in this paper are as follows. In a nutshell, we tighten the frontier between the tractable and intractable cases of the Telephone Broadcasting problem.
  • We prove that Telephone Broadcasting is NP-Complete in two of the simplest sparse graph families: $k$-cycle graphs (Theorem 3.4) and $k$-path graphs (Theorem 3.5). Both families have been studied in the broadcasting literature, yet their complexity status remained open. Our hardness results also imply hardness for broader classes, including series–parallel graphs (SP-graphs), for which no hardness was previously known.
  • We design polynomial-time approximation schemes (PTAS) for both $k$-cycle graphs (Corollary 4.7) and $k$-path graphs (Corollary 4.14). To the best of our knowledge, these are the first graph families for which Telephone Broadcasting admits both a hardness result and a matching PTAS, highlighting their role as boundary cases between tractability and intractability.
  • We present a polynomial-time algorithm for optimal broadcasting in graphs of bounded bandwidth (Theorem 5.2). This result generalizes several known tractable cases, such as chains of rings and related structures. Notably, Telephone Broadcasting is among the few problems exhibiting a complexity gap between graphs of bounded bandwidth (where the problem is tractable) and graphs of bounded pathwidth (where the problem is NP-hard). In particular, $k$-cycle and $k$-path graphs both have pathwidth 2 but unbounded bandwidth, placing them precisely on the intractable side of this gap.
(a) A snowflake graph (cactus of pathwdith 2) [1]
(b) Graph of distance 1 to path forest [6]
(c) A $k$-cycle graph (Theorem 3.4)
(d) A $k$-path graph (Theorem 3.5)
Figure 1: Sparse graph families for which telephone broadcasting is proven to be NP-hard.
1.2 Overview of Techniques
Our results combine hardness reductions, approximation schemes, and dynamic programming, each tailored to the structural properties of the graph families under consideration. We give a brief outline here.
  • For the negative results, we design reductions that encode scheduling constraints directly into the graph structure. In particular, the NP-hardness proofs for $k$-cycle and $k$-path graphs exploit the interaction between cycle (or path) lengths and the timing of broadcast calls. The shared vertex in $k$-cycles (or the pair of endpoints in $k$-paths) acts as the originator, and we show that deciding feasibility becomes equivalent to a combinatorial partitioning problem whose complexity propagates to the broadcasting instance. Our reductions build a more restricted variant of Numerical 3-Dimensional Matching (RN3DM) (Definition 3.1), previously employed in hardness proofs for telephone broadcasting. This restricted variant, which is known to be still NP-hard [22], makes it possible to encode number-theoretic constraints using only these simple graph families.
  • We design polynomial-time approximation schemes for $k$-cycle and $k$-path graphs by reducing the broadcasting problem to covering problems over integers. For $k$-cycle graphs, the task reduces to covering a multiset of cycle lengths $\{\ell_{1},\ldots,\ell_{k}\}$ with the smallest possible broadcast round $\beta$ using pairs of numbers from $\beta\,]$, so that for each $\ell_{i}$ there exist $a,b\in[\,\beta\,]$ associated with $\ell_{i}$ such that $a+b\geq\ell_{i}$ (See Definition 4.1). For $k$-path graphs, the covering problem is slightly different: we seek the smallest $\beta$ such that each $\ell_{i}$ can be covered by one number from $\beta\,]$ and one from $\beta-d\,]$, where $d$ is the length of the shortest path (see Definition 4.9). In both cases, approximate solutions to the number-theoretic covering problems translate into near-optimal broadcasting schedules, yielding PTASs for $k$-cycle and $k$-path graphs.
  • For finding optimal broadcast schemes in graphs of bounded bandwidth, we exploit the fact that a bandwidth-$k$ ordering restricts each vertex to neighbors within a sliding window of $\mathcal{O}(k)$ vertices. This allows us to describe the status of any partial broadcast tree using a compact boundary state and use a dynamic programming approach which has a table with $n$ rows (one for each prefix of the ordering) and at most $n^{\mathcal{O}(k^{2})}$ states per row, giving a total size bounded by $f(k)\cdot n^{\mathcal{O}(k^{2})}$.
2 Preliminaries
We use notation $[n]$ and $[n,m]$ for positive integers $n$ and $m$ for the set of all natural numbers less than or equal to $n$ and all natural numbers between $n$ and $m$ (inclusive), respectively (i.e. $n]=\{1,2,\ldots,n\}$ and $n,m]=\{n,n+1,\ldots,m\}$). Also, we refer to the number of distinct elements in a multiset as its “support”.
Definition 2.1 (Telephone Broadcasting [16]).
An instance $(G,s)$ of the Telephone Broadcasting problem is defined by a connected, undirected, and unweighted graph $G=(V,E)$ and a vertex $s\in V$, where $s$ is the only informed vertex. The broadcasting protocol is synchronous and occurs in discrete rounds. In each round, an informed vertex can inform at most one of its uninformed neighbors. The goal is to broadcast the message as quickly as possible so that all vertices in $V$ get informed in the minimum number of rounds.
Given that the problem can be solved in linear time on trees [10], it is common to represent a broadcast scheme—that is, a feasible solution to the problem—by a spanning tree of $G$ rooted at $s$. We refer to such a tree as a broadcast tree. In this tree, an edge $(u,v)$ in the tree indicates that vertex $v$ receives the message from $u$ in the broadcast process. Thus, throughout the paper, the terms “broadcast tree” and “broadcast scheme” are used interchangeably. Likewise, a broadcast subtree denotes a partial solution, i.e., a subtree of a broadcast tree capturing how a subset of vertices is informed. Throughout, we use ${\boldsymbol{br}^{*}}(G,s)$ to denote the optimal broadcast time of graph $G$ when the originator is $s$. When the context is clear, we simply write $\boldsymbol{br}^{*}$} in place of ${\boldsymbol{br}^{*}}(G,s)$.
Definition 2.2 ($k$-cycle and $k$-path graphs).
A $k$-cycle graph $G$ contains $k$ cycles with exactly one common vertex $r$. A $k$-path graph $G$ contains $k$ paths with shared endpoints $s$ and $t$.
See Figure 1 (parts (c) and (d)) for illustrations of $k$-cycle and $k$-path graphs. Observe that any $k$-cycle (respectively, $k$-path) graph can be fully described by a set of integers specifying the number of vertices on each cycle (respectively, path). This representation allows us to relate such graphs directly to integer sets, a connection that underlies both our hardness proofs and our algorithmic techniques. These graph families are also simple representatives of well-studied classes: both have bandwidth 2, $k$-cycle graphs form a subclass of cactus graphs, and $k$-path graphs form a subclass of series–parallel graphs. Consequently, our hardness results extend beyond $k$-cycles and $k$-paths to broader sparse families, including series–parallel graphs, for which no hardness results were previously known.
Definition 2.3 (Graph Bandwidth).
Given a graph $G$, an ordering with bandwidth $k$ of $G$ is a permutation $v_{1},\ldots,v_{n}$ of vertices of $G$ such that for each edge $v_{i}\sim v_{j}$ we have $|i-j|\leq k$. The bandwidth of $G$ is the minimum $k$ such that such an ordering exists.
Certain graphs of bounded bandwidth have also been studied in the context of telephone broadcasting. For instance, necklace graphs—collections of cycles where each consecutive pair shares a vertex—have bandwidth $k=2$ and the problem can be optimally solved for them [13, 15].
3 Hardness Results
In this section, we present our hardness results for Telephone Broadcasting in $k$-cycle graphs and $k$-path graphs. These results extend the existing hardness landscape by showing that even some of the simplest graph families exhibit computational intractability.
Recent hardness results for broadcasting typically rely on reductions from variants of the Numerical Matching with Target Sums (NMTS) problem, which is known to be strongly NP-hard [11] (Problem SP17). An instance of NMTS consists of three multisets $U,V,$ and $W$, each containing $m$ positive integers. The decision problem asks whether there exists a partition of $U\cup V\cup W$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, such that each $A_{i}$ contains exactly one element $u_{i}\in U$, one element $v_{i}\in V$, and one element $w_{i}\in W$, with the property that $u_{i}+v_{i}=w_{i}$ for all $i\in[m]$.
Tale [21] introduced a variant of NMTS, called Numerical 3-Dimensional (Almost) Matching, in which no repeated entries appear in $U$ and $V$, and all elements of $U\cup V\cup W$ are multiples of $m$ and two integers $T$ and $V$. As in NMTS, the objective is to partition $U\cup V$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, each containing exactly one element from $U$ and one element from $V$. However, instead of requiring $\sum_{a\in A_{i}}a=w_{i}$, the condition is relaxed to $\sum_{a\in A_{i}}a\geq T-\lambda.$ Tale showed that this problem is strongly NP-hard via a reduction from NMTS, and subsequently reduced it to Telephone Broadcasting in graphs of pathwidth at most 3. Egami et al. [6] studied another restricted variant of NMTS, in which all members of $U$ are even while all members of $V\cup W$ are odd. They proved that this variant remains strongly NP-hard and provided a reduction to broadcasting in graphs that are distance-1 from a path forest.
To establish our hardness results, we rely on another restricted version of NMTS in which $U=V=[m]$. This problem, referred to as Restricted Numerical 3-Dimensional Matching (RN3DM), was studied by Yu et al. [22], who established its strong NP-hardness. Their proof relies on a reduction from the two-machine flowshop scheduling problem with delays, which they first showed to be NP-hard. Formally, RN3DM is defined as follows.
Definition 3.1 (Restricted Numerical 3-Dimensional Matching (RN3DM) [22]).
Given a multiset $W=\{w_{1},w_{2},\ldots,w_{m}\}$ of positive integers such that $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$, do there exist two permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)+w_{i}=e\quad\text{for all }1\leq i\leq m\,$?
Example 1 (Yes-instance of RN3DM).
Consider the instance of RN3DM with $W=\{1,3,4,4\}$, so $m=4$. Note that we have $\sum_{i=1}^{m}w_{i}+m(m+1)=32=me$ for $e=8$. Let $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$. One can check that for every $i\in[4]$, $\lambda(i)+\mu(i)+w_{i}=e.$ For instance, $\lambda(1)+\mu(1)+w_{1}=4+3+1=8$, and $\lambda(2)+\mu(2)+w_{2}=1+4+3=8.$ The same holds for $i=3,4$. Hence $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
Given that RN3DM is strongly NP-hard [22], we assume that the numbers in $W$ are encoded in unary. We present reductions from RN3DM to broadcasting in $k$-cycle and $k$-path graphs. For $k$-cycle graphs, our approach introduces an intermediate problem as part of the reduction series, whereas for $k$-path graphs, we directly reduce RN3DM to telephone broadcasting.
3.1 Hardness of Telephone Broadcasting in $k$-cycle graphs
We prove that the $k$-cycle broadcasting decision problem is NP-hard through a series of reductions starting from RN3DM. As a first step, we define an intermediate problem, called Even-Odd RN3DM, which is a variant of RN3DM. In this variant, instead of using two permutations $\lambda$ and $\mu$ of $[m]$ (matched with the input set $W$), we use $\alpha$ as a permutation of the even integers in $[2m]$ and $\beta$ as a permutation of the odd integers in $[2m]$ to be matched with an input set that we call $C$. Formally, Even-Odd RN3DM is defined as follows.
Definition 3.2 (Even-Odd Restricted Numerical 3-Dimensional Matching (Even-Odd RN3DM)).
Given a multiset $C=\{c_{1},c_{2},\ldots,c_{m}\}$ of positive integers such that $\sum_{i=1}^{m}c_{i}=m(2m+1),$ does there exist a permutation $\alpha$ of $2,4,6,\ldots,2m$} and a permutation $\beta$ of $1,3,5,\ldots,2m-1$} such that $\alpha(i)+\beta(i)=c_{i}\quad\text{for all }1\leq i\leq m\,?$
Observe that any instance of Even-Odd RN3DM in which $C$ contains an even element is trivially a No-instance, since each $c_{i}$ must be represented as the sum of one even and one odd number, which always yields an odd value. Hence, without loss of generality, we restrict attention to instances where all elements of $C$ are odd.
Example 2 (Yes-instance of Even-Odd RN3DM).
Consider the instance of Even-Odd RN3DM with $C=\{7,7,9,13\}$, where $m=4$. Note that $\sum_{i=1}^{4}c_{i}=36=m(2m+1)$. Now take the permutations $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$. One can check for every $i\in[4]$, $\alpha(i)+\beta(i)=c_{i}$. For instance $\alpha(1)+\beta(1)=4+3=7=c_{1}$, and $\alpha(2)+\beta(2)=6+1=7=c_{2}$. The same holds for $i=3,4$. Thus, $\alpha$ and $\beta$ certify that $C$ is a yes-instance of Even-Odd RN3DM.
[…]
Figure 2: An instance $C=\{7,7,9,13\}$ of Even-Odd RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=8$). The broadcast tree associated with with solution $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$ is highlighted.
Lemma 3.3.
Even-Odd RN3DM is strongly NP-hard.
Proof Sketch: We use a reduction from RN3DM, which is strongly NP-Hard [22]. Given an instance $W=\{w_{1},\ldots,w_{m}\}$ of RN3DM with target sum $e$, we construct an instance $C=\{c_{1},\ldots,c_{m}\}$ of Even-Odd RN3DM by setting $c_{i}=2e-2w_{i}-1\quad\text{for each }i\in[m].$ It is straightforward to verify that $W$ is a Yes-instance of RN3DM if and only if $C$ is a Yes-instance of Even-Odd RN3DM. The transformation uses unary encodings, so the reduction is strong. $\Box$
Proof. We give a polynomial-time reduction from RN3DM. Let $W=\{w_{1},\ldots,w_{m}\}$ be an arbitrary instance of RN3DM. By definition, we have $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$. Construct an instance $C=\{c_{1},\ldots,c_{m}\}$ of Even-Odd RN3DM, where $c_{i}=2e-2w_{i}-1\quad\text{for each }i\in[m].$ To illustrate, the instance of RN3DM in Example 1 reduces to the instance of Even-Odd RN3DM in Example 2. First, we verify that $C$ satisfies the side condition of Even-Odd RN3DM:
$\sum_{i=1}^{m}c_{i}=\sum_{i=1}^{m}(2e-2w_{i}-1)=2em-2(\sum_{i=1}^{m}w_{i})-m=2em-2(me-m(m+1))-m=m(2m+1).$
In the above equalities, we used the definition of $C$ and the side condition that holds for the instance $W$ of RN3DM. Thus, $C$ is a valid instance of Even-Odd RN3DM.
Next, we show the correctness of the reduction.
(Forward direction). Suppose $W$ is a yes-instance of RN3DM. Then there exist permutations $\lambda$ and $\mu$ of $[m]$ such that for each $i\in[m]$, $\lambda(i)+\mu(i)+w_{i}=e.$ Define $\alpha$ as the permutation of $\{2,4,\ldots,2m\}$ given by $\alpha(i)=2\lambda(i)$, and $\beta$ as the permutation of $\{1,3,\ldots,2m-1\}$ given by $\beta(i)=2\mu(i)-1$. Then, for every $i$,
$\alpha(i)+\beta(i)=2\lambda(i)+(2\mu(i)-1)=2(\lambda(i)+\mu(i))-1=2(e-w_{i})-1=2e-2w_{i}-1=c_{i}.$
Hence, $\alpha$ and $\beta$ certify that $C$ is a yes-instance of Even-Odd RN3DM.
(Reverse direction). Suppose $C$ is a yes-instance of Even-Odd RN3DM. Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that for each $i\in[m]$, $\alpha(i)+\beta(i)=c_{i}.$ Define $\lambda(i)=\alpha(i)/2$ and $\mu(i)=(\beta(i)+1)/2$. Both $\lambda$ and $\mu$ are permutations of $[m]$. For each $i$,
$\lambda(i)+\mu(i)+w_{i}=\frac{\alpha(i)}{2}+\frac{\beta(i)+1}{2}+w_{i}=\frac{c_{i}+1}{2}+w_{i}=\frac{(2e-2w_{i}-1)+1}{2}+w_{i}=e.$
Thus $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
Since the reduction preserves yes-instances in both directions and can be carried out in polynomial time, Even-Odd RN3DM is strongly NP-hard. $\Box$
Theorem 3.4.
Broadcasting in $k$-cycle graphs is NP-complete.
Proof. Telephone Broadcasting is in NP in general graphs (and hence also for $k$-cycle graphs). We prove NP-hardness by a reduction from the Even-Odd RN3DM problem, which is strongly NP-Hard by Lemma 3.3.
Let $C=\{c_{1},c_{2},\ldots,c_{m}\}$ be an instance of Even-Odd RN3DM. We construct an instance $(G,s,t)$ of Telephone Broadcasting in $k$-cycle graphs as follows. Graph $G$ consists of $m$ cycles, containing $c_{1},c_{2},\ldots,c_{m}$, extra vertices in addition to a single vertex $s$ that is shared between them. Now $s$ serves as the originator of the broadcast. Thus, the number of vertices in $G$ is $n=\sum_{i=1}^{m}c_{i}+1=m(2m+1)+1,$ and the degree of $s$ is $2m$. Since RN3DM is strongly NP-hard by Lemma 3.3, we may assume the values $c_{i}$ are encoded in unary, so the reduction is polynomial time. The decision problem asks whether broadcasting completes within $t=2m$ rounds. See Figure 2.
(Forward direction). Suppose the answer to instance $C$ of Even-Odd RN3DM is yes. Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that $\alpha(i)+\beta(i)=c_{i}$ for all $i\in[m]$. Construct a broadcast scheme as follows: the source $s$ informs its two neighbors in each cycle $c_{i}$ at times $2m-\alpha(i)+1$ and $2m-\beta(i)+1$, respectively. Given that $\alpha$ and $\beta$ are permutations of space $[2m]$, the source $s$ informs exactly one vertex at each given round $i\in[2m]$. Each newly informed vertex continues by informing its uninformed neighbor in the subsequent round. Hence, by round $br=2m$, exactly $\alpha(i)+\beta(i)=c_{i}$ vertices of the cycle (excluding $s$) are informed, and therefore the entire graph is informed. Thus, $(G,s,br)$ is a yes-instance of broadcasting.
(Reverse direction). Now suppose $(G,s,br)$ is a yes-instance of broadcasting with $br=2m$. So, there is a broadcast scheme that completes within $22m$ rounds. Without loss of generality, assume the scheme is non-lazy, i.e., every informed vertex calls an uninformed neighbor whenever possible. Let $t_{i}$ denote the number of vertices newly informed in round $i$. At round 0, only $s$ is informed: $t_{0}=1$. At each round $i$, only $s$ and the vertices informed at time i1i-1 can inform new vertices (since other vertices of degree 2 have no additional uninformed neighbors left). Hence, $t_{i+1}\leq t_{i}+1$. It follows that the number of informed vertices by round $i$ is at most $1+1+2+\cdots+i=1+\frac{i(i+1)}{2}$. For $i=2m$, this reaches exactly $1+m(2m+1)=n$, so the broadcast must finish in exactly $2m$ rounds (and not earlier). This forces the source $s$ to make a call in every round, and since $\deg(s)=2m$, $s$ informs exactly one neighbor per round. Consequently, each cycle $c_{i}$ must be partitioned into two paths, informed from $s$ in opposite directions, of lengths $\alpha(i)$ and $\beta(i)$ with $\alpha(i)+\beta(i)=c_{i}$. Because $c_{i}$ is odd (otherwise the Even-Odd RN3DM instance would be trivial), we conclude that one of $\alpha(i),\beta(i)$ is odd and the other is even. This provides the exact decomposition required for a yes-instance of Even-Odd RN3DM. $\Box$
3.2 Hardness of Telephone Broadcasting in $k$-Path Graphs
Next, we use a reduction from RN3DM to establish NP-completeness of broadcasting in $k$-path graphs.
Theorem 3.5.
Broadcasting in $k$-path graphs is NP-complete.
Proof. Telephone Broadcasting belongs to NP for general graphs and hence also for $k$-path graphs. We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-Hard [22].
Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM. By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$. We construct an instance $(G,s,br)$ of Telephone Broadcasting in $k$-path graphs as follows. The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$. Furthermore, we add an edge connecting $s$ and $t$. The vertex $s$ serves as the source of the broadcast. The total number of vertices in $G$ is $n=2+\sum_{i=1}^{m}(e-w_{i})=2+me-\sum_{i=1}^{m}w_{i}=2+m(m+1).$ Both $s$ and $t$ have degree $2m+1$. Since RN3DM is strongly NP-hard, we may assume that the values $w_{i}$ are encoded in unary, ensuring that the reduction is polynomial time. The decision problem asks whether broadcasting completes within $br=m+1$ rounds (see Figure 3).
(Forward direction). Suppose $W$ is a yes-instance of RN3DM. Then there exist permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)=e-w_{i}=\ell_{i}$ for all $i\in[m]$. Construct a broadcast schedule as follows: in round 1, the source $s$ calls $t$. Subsequently, in round $j\in[m]$, $s$ and $t$ inform their neighbors in the $j$’th path at times $m+1-\lambda(j)$ and $m+1-\mu(j)$, respectively. Since $\lambda$ and $\mu$ are permutations, exactly one new vertex is informed from $s$ and one from $t$ in each round (i.e., this is a valid broadcast scheme). Each newly informed vertex continues the broadcast by informing its remaining neighbor in the next round. Hence, by round $br=m+1$, exactly $\lambda(j)$ vertices from the side of $s$ and $\mu(j)$ vertices from the side of $t$ from path $j$ are informed; therefore, all $\ell_{j}$ internal vertices of every path are informed, and the entire graph is covered. Thus, $(G,s,br)$ is a yes-instance of Telephone Broadcasting.
(Reverse direction). Suppose $(G,s,br)$ is a yes-instance of Telephone Broadcasting with $br=m+1$. Without loss of generality, assume the scheme is non-lazy, i.e., every informed vertex calls an uninformed neighbor whenever possible. Let $t_{i}$ denote the number of vertices newly informed in round $i$. At round 0, only $s$ is informed, so $t_{0}=1$. Suppose $t$ is informed at round $p\geq 1$. At each round $i$, only $s$, possibly $t$ (if $i>p$), and the vertices informed at round i1i-1 can make new calls. Thus $t_{i+1}\leq t_{i}+1$ for $i\leq p$, and $t_{i+1}\leq t_{i}+2$ for $i\geq p$. This upper bound value is maximized when $p=1$, i.e., when $s$ calls $t$ immediately. In that case, we have $t_{0}=1$, $t_{1}=1$, $t_{2}=2$, and for $i\geq 2$, $t_{i}=t_{i-1}+2$. Therefore, the total number of informed vertices by round $i$ is at most $1+1+2+4+6+\cdots+2(i-1)=i(i-1)+2.$ For $i=m+1$, this reaches exactly $n=2+m(m+1)$. Hence, the broadcast cannot finish in less than $m+1$ rounds, i.e., must finish in exactly $m+1$ rounds. This forces $s$ to call $t$ in round 1, and both $s$ and $t$ to call exactly one new neighbor in each subsequent round. Otherwise, the number of informed vertices at time $t_{i}$ will be less than the above upper bound, and hence the broadcast cannot complete within $m+1$ rounds.
Consequently, each path of length $\ell_{i}$ must be partitioned into two subpaths, one informed from $s$ and the other from $t$, of lengths $\lambda(i)$ and $\mu(i)$ such that $\lambda(i)+\mu(i)=\ell_{i}$. Recall that $\ell_{i}=e-w_{i}$. Since both $s$ and $t$ inform exactly one new neighbor in each round, it follows that for any two distinct paths $i$ and $j$, we must have $\lambda(i)\neq\lambda(j)$ and $\mu(i)\neq\mu(j)$. In other words, $\lambda$ and $\mu$ are permutations of $[m]$. This corresponds exactly to a yes-instance of RN3DM. $\Box$
[…]
Figure 3: An instance $W=\{1,3,4,4\}$ of RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=5$). The broadcast tree associated with with solution $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$ is highlighted.
4 Polynomial Time Approximation Schemes
4.1 PTAS for $k$-Cycle Graphs
In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-cycle graphs. Recall that a $k$-cycle graph is obtained by joining $k$ cycles at a common central vertex $r$. We assume that the number of cycles kk—and hence the optimal broadcast time—is asymptotically large; for constant $k$, the problem can be solved by a straightforward exhaustive search (see, e.g., [4]).
In what follows, we design a PTAS for the case where the broadcast originates at $s=r$. This restriction is without loss of generality: any $c$-approximation for the case that the center is the originator (when $s=r$) can be extended to a $c$-approximation in the general case ($s\neq r$). Suppose the originator $s$ is a non-center vertex on some cycle $C_{m}$, at distance $d\geq 1$ from $r$. The message can first be forwarded along the shortest path from $s$ to $r$. Once $r$ is informed, it first sends the message to its other neighbor on $C_{m}$ and then proceeds with the $c$-approximation algorithm on the rest of the graph. If the overall broadcast time is determined by a vertex in $C_{m}$, then the scheme is optimal. Otherwise, the process completes within $d+1+c\cdot\boldsymbol{br}^{-}$ rounds, where $\boldsymbol{br}^{-}$ is the optimal broadcast time excluding $C_{m}$. Since ${\boldsymbol{br}^{*}}\geq d+\boldsymbol{br}^{-}$, this yields the same approximation factor $c$.
We introduce the following problem, which is the key part of our PTAS algorithms.
Definition 4.1.
The input to the Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers. The objective is to determine the minimum integer $m$ such that the set $C=[m]$ can be partitioned into $n$ subsets $C_{0},C_{1},\ldots,C_{n}$, such that for each $i\in[n]$, we have $|C_{i}|\leq 2$ and $\sum_{x\in C_{i}}x\,\geq\,\ell_{i}$. In this context, we say that $\ell_{i}$ is covered by the elements of $C_{i}$, and we refer to $C$ as the covering set of $S$.
Example 3.
When $S=\{13,8,7,6\}$, the optimal solution is given by $C=[8]$, and we have $C_{1}=\{8,5\},C_{2}=\{7,2\},C_{3}=\{4,3\},C_{4}=\{6\}$ (and $C_{0}=\{1\}$).
We begin by establishing the equivalence between Telephone Broadcasting in $k$-cycle graphs and the Prefix Covering problem.
Lemma 4.2.
Let $G$ be a $k$-cycle graph with central vertex $s$, where the $i$-th cycle contains $c_{i}$ additional vertices besides $s$, and let $S=\{c_{1},\ldots,c_{k}\}$. Then Telephone Broadcasting for instance $(G,s)$ becomes equivalent to Prefix Covering for $S$.
Proof Sketch: Let $\boldsymbol{br}$ denote the number of broadcast rounds. If the center $s$ informs two neighbors of cycle $i$ at rounds $x$ and $y$, then the remaining $c_{i}$ vertices are informed within $\boldsymbol{br}$ rounds if and only if we have $(\boldsymbol{br}-x+1)+(\boldsymbol{br}-y+1)\geq c_{i}$. This is equivalent to covering $c_{i}$ by two elements of $\boldsymbol{br}\,]$ whose sum is at least $c_{i}$. Thus, a broadcast schedule with $\boldsymbol{br}$ rounds corresponds to a feasible Prefix Covering solution, and conversely any such covering yields a valid broadcast schedule. $\Box$
Proof. We show that any feasible broadcasting scheme with $\boldsymbol{br}$ rounds, for instance $(G,s)$, corresponds to a feasible covering of $S$ using $\boldsymbol{br}]$, and conversely any such covering yields a valid broadcasting scheme. This one-to-one correspondence ensures that the two problems have the same optimal solutions.
From Telephone Broadcasting to Prefix Covering: In any broadcast scheme with $\boldsymbol{br}$ rounds, $r(=s)$ must call at most two vertices on each cycle. If $r$ informs neighbors in cycle $i$ at rounds $x$ and $y$, then the remaining $c_{i}-1$ vertices of cycle $i$ must be covered within the remaining rounds, which is feasible exactly when $(\boldsymbol{br}-x+1)+(\boldsymbol{br}-y+1)\geq c_{i}$. Thus, each $c_{i}$ is covered by the pair $\{\boldsymbol{br}-x+1,\boldsymbol{br}-y+1\}$ in the set $\boldsymbol{br}]$. Since $r$ uses distinct rounds for distinct neighbors, the covering also respects the injectivity requirement. Hence, every broadcast scheme with $\boldsymbol{br}$ rounds yields a valid covering for $S$ with $\boldsymbol{br}]$.
From Prefix Covering to Telephone Broadcasting: Conversely, suppose $S$ has a covering using $C=[\boldsymbol{br}]$. If $c_{i}$ is covered by $x$ and $y$, then in the broadcast scheme $r$ informs two neighbors in the $i$’th cycle at rounds $\boldsymbol{br}-x+1$ and $\boldsymbol{br}-y+1$. Because $x+y\geq c_{i}$, the message traverses all of $i$’th cycle within $\boldsymbol{br}$ rounds. Injectivity of the covering ensures that $r$ never attempts more than one call in a round. Thus, every covering for the Prefix Covering instance corresponds to a valid broadcast scheme with $\boldsymbol{br}$ rounds.
Therefore, Telephone Broadcasting in $k$-cycle graphs is equivalent to the Prefix Covering problem. $\Box$
4.1.1 PTAS for Prefix Covering
In this section, we present a PTAS for the Prefix Covering problem, that is, an algorithm with approximation factor $(1+\epsilon)$ for a given $\epsilon>1$. By Lemma 4.2, this would give a PTAS for Telephone Broadcasting in $k$-cycle graphs. As a high-level view of the algorithm, it addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the covering set into a simplified version with constant support (distinct elements), making it feasible to apply an exhaustive search which maintains provable approximation guarantees.
The algorithm operates as follows. Given an instance $S$ of Prefix Covering, we first construct a rounded set $\mathfrak{R}_{p}(S)$, where each element of $S$ is rounded up to a larger value in $S$ so that the support of $\mathfrak{R}_{p}(S)$ has size $p$. Here $p$ is a constant integer chosen as a function of $\epsilon$. Since rounding increases item values, covering $\mathfrak{R}_{p}(S)$ may require a larger covering set. Nevertheless, we show that the optimal covering of $\mathfrak{R}_{p}(S)$ is within a factor $(1+2/p)$ of the optimal covering of $S$ (Proposition 4.4). Thus, it suffices to focus on covering $\mathfrak{R}_{p}(S)$ rather than $S$. For that, we use a binary search over candidate values $m$, where $m$ specifies the size of the covering set $C=[m]$. For each $m$, we define a rounded version $\mathfrak{R}_{p}(C)$, obtained by rounding up the elements of $C$ so that its support also has size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$, then it can also be covered using $C^{\prime}=[m^{\prime}]$ with $m^{\prime}\leq(1+1/p)m$ (4.1). Conversely, if $\mathfrak{R}_{p}(S)$ cannot be covered by $\mathfrak{R}_{p}(C)$, then it cannot be covered by $C(=[m])$. Hence, verifying whether $\mathfrak{R}_{p}(C)$ successfully covers $\mathfrak{R}_{p}(S)$ is sufficient for approximating the original instance. This verification can be performed by exhaustive search over $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$. Since both sets have constant support (bounded by $p$), this step can be carried out in time independent of $n$ (Lemma 4.5). In what follows, we provide the details of the algorithm.
Based on Definition 4.1, an instance of Prefix Covering can be defined as a tuple $(S,C)$, where $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ and $C=[m]$ are the ground multiset and the covering set, respectively. Throughout this section, we assume that $\ell_{1}\geq\ell_{2}\geq\ldots\geq\ell_{n}$.
Definition 4.3.
Let $S$ be a finite set and let $p$ be a positive integer. The rounded set $\mathfrak{R}_{p}(S)$ of $S$ with $p$ parts is constructed as follows:
  1. (i).
    Order the elements of $S$ in non-increasing sequence: $s_{1}\geq s_{2}\geq\cdots\geq s_{|S|}$.
  2. (ii).
    Partition this ordered sequence into $p$ consecutive parts $P_{1},P_{2},\ldots,P_{p}$, where:
    • For $i=1,2,\ldots,p-1$: $P_{i}$ contains exactly $\lceil|S|/p\rceil$ elements,
    • $P_{p}$ contains the remaining $|S|-(p-1)\lceil|S|/p\rceil$ elements.
  3. (iii).
    For each part $P_{i}$, let $m_{i}$ denote the maximum element in $P_{i}$. Then $\mathfrak{R}_{p}(S)$ is the multiset obtained by replacing each element in $P_{i}$ with $m_{i}$ for all $i\in\{1,2,\ldots,p\}$.
Example 4.
For $S=\{123,67,65,45,43,43,43,18,12,12,10,6,4,4,1,1\}$ and $p=4$, we get $\mathfrak{R}_{4}(S)=\{123,123,123,123,43,43,43,43,12,12,12,12,4,4,4,4\}$.
Proposition 4.4.
$\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq\boldsymbol{OPT}(S)+2\lceil|S|/p\rceil$, where $\boldsymbol{OPT}$ is the optimal cover for a specific instance of Prefix Covering.
Proof. By definition, $\mathfrak{R}_{p}(S)$ consists of $p$ parts. Given an optimal cover of $S$, form a cover of $\mathfrak{R}_{p}(S)$ by assigning the numbers given to the part $i$ in $S$ to the part $(i+1)$ in $\mathfrak{R}_{p}(S)$ for $i\in[p-1]$. As the size and value of part $i+1$ is not greater than the value of part $i$, the set of numbers that can cover part $i$ can also cover part $i+1$.
Therefore, the first part is the only part that needs to be covered. We show that adding an extra $2\lceil|S|/p\rceil$ to the covering set is enough to cover the first part. This is because every $\ell_{i}\in S$ is less than $2$ items of size $\boldsymbol{OPT}(S)$, as otherwise it would not be possible to cover $\ell_{i}$ with at most two numbers in $\boldsymbol{OPT}(S)]$. Thus, by adding an extra $2\lceil|S|/p\rceil$ items to the covering set, we can cover $\mathfrak{R}_{p}(S)$. $\Box$
Example 5 (Based on Example 4).
In Example 4, items in a covering set that were used to cover $\{123,67,65,45\}$ in the optimal cover of $S$ will be used to cover $\{43,43,43,18\}$ in $\mathfrak{R}_{p}(S)$; those used to cover $\{43,43,43,18\}$ will be used to cover $\{12,12,12,12\}$ in $\mathfrak{R}_{p}(S)$, and so on. The only remaining uncovered part in $\mathfrak{R}_{p}(S)$ will be the first part ($(123,123,123,123)$), which has size $|S|/p$. Moreover, every element in that part has size at most $2\boldsymbol{OPT}(S)$. Therefore, the first part in $\mathfrak{R}_{p}(S)$ can be covered by an extra $2|S|/p$ in the covering set. In the example, if $\boldsymbol{OPT}(S)=73$ in the above example, we can cover $\{123,123,123,123\}$ by adding $\{74,75,76,77,78,79,80,81\}$ to the optimal cover set of $S$ to cover $\mathfrak{R}_{p}(S)$ using $\boldsymbol{OPT}(S)+2|S|/p$ rounds.
Given that $\boldsymbol{OPT}(S)\geq|S|$, we can conclude that an optimal covering of $\mathfrak{R}_{p}(S)$ gives a ($1+2/p$) approximation for covering $S$. Therefore, we can now focus on covering $\mathfrak{R}_{p}(S)$, which has constant support $p$.
Let $m$ be a given value and let $C=[m]$. We present a procedure that either constructs a covering of $\mathfrak{R}_{p}(S)$ using the set $C^{\prime}=[m^{\prime}]$, where $m^{\prime}=m+\lceil m/p\rceil$, or certifies that $\mathfrak{R}_{p}(S)$ cannot be covered using $[m]$. For this purpose, we use the rounded set $\mathfrak{R}_{p}(C)$ to cover $\mathfrak{R}_{p}(S)$. We define the stretched covering $\mathfrak{R}_{p}(C)$ as the set obtained by partitioning the $C$ and rounding up its elements (as in Definition 4.3), ensuring that the covering set has constant support $p$. This enables a binary-search strategy to find the smallest feasible value of $m$.
Example 6.
Suppose $\mathfrak{R}_{p}(S)=\{123,123,123,123,43,43,43,43,12,12,12,12,4,4,4,4\}$ and $p=3$. For $m=12$, we have $C=\{12,11,10,9,8,7,6,5,4,3,2,1\}$, and we aim to cover $\mathfrak{R}_{p}(S)$ with $\mathfrak{R}_{p}(C)=\{12,12,12,12,8,8,8,8,4,4,4,4\}$.
Claim 4.1.
Given that we can cover $S$ with $\mathfrak{R}_{p}(C)$, where $C=[m]$, we can cover $S$ using $C^{\prime}=[m^{\prime}]$, where $m^{\prime}=m+\lceil m/p\rceil$.
Proof. Based on definition, $\mathfrak{R}_{p}(C)$ contains $p$ parts. Suppose we have a covering with $\mathfrak{R}_{p}(C)$ for $S$. Associate value $m+\lceil m/p\rceil-(i-1)$ to the $i$’th largest number in $C$ for $i\in[m]$. This transformation only increases the value as the first block with original value $m$ matches to $m+1,\ldots,m+\lceil m/p\rceil$, the second block matches to $m-\lceil m/p\rceil+1,\ldots,m$, and so on. More specifically, the $i$’th block matches to $m-(i-1)\lceil m/p\rceil+1,\ldots,m-(i-2)\lceil m/p\rceil$, which is greater than the original value of the block. Now, as the new set is a subset of $C^{\prime}$ and every value in $C$ is larger than its original value, the same covering can cover $S$ with the corresponding values in $C^{\prime}$. $\Box$
Example 7 (Based on Example 6).
If we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)=\{12,12,12,12,8,$ 8,8,8,4,4,4,4}8,8,8,4,4,4,4\}, we can cover $\mathfrak{R}_{p}(S)$ using $C^{\prime}=\{16,15,14,13,12,11,10,9,8,7,6,5,4,3,2,1\}$: items of $\mathfrak{R}_{p}(S)$ that were covered by $(12,12,12,12)$ in $\mathfrak{R}_{p}(C)$ can be covered with $(16,15,14,13)$ in $C^{\prime}$, whatever was covered by $(8,8,8,8)$ in $\mathfrak{R}_{p}(C)$, can be covered by $(12,11,10,9)$ in $C^{\prime}$, etc. Hence, if we can cover $\mathfrak{R}_{p}(S)$ with $\mathfrak{R}_{p}(C)$, we can cover $\mathfrak{R}_{p}(S)$ using $C^{\prime}$; which means we can cover $\mathfrak{R}_{p}(S)$ using $m^{\prime}]$.
Lemma 4.5.
Given two sets $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering set $\mathfrak{R}_{p}(C)$ (if one exists) or correctly reports that no such covering is possible.
Proof. Both the ground set $\mathfrak{R}_{p}(S)$ and the covering set $\mathfrak{R}_{p}(C)$ have constant support $p$. Fix a distinct value $v\in\mathfrak{R}_{p}(S)$. For each pair $(a,b)$ of elements from $\mathfrak{R}_{p}(C)$, we can encode how many times $(a,b)$ is used to cover elements of $\mathfrak{R}_{p}(S)$ equal to $v$. Since there are at most $|S|$ such elements, this count can be represented in $\log|S|$ bits. As there are at most $p^{2}$ pairs from $\mathfrak{R}_{p}(C)$ and $p$ distinct values in $\mathfrak{R}_{p}(S)$, any candidate covering can be encoded using at most $p^{3}\log|S|$ bits. Hence, the number of candidate solutions is bounded by $\mathcal{O}(2^{p^{3}}|S|)$. Checking whether a candidate encoding yields a valid covering takes $\mathcal{O}(|S|+|C|)$ time. Therefore, the total running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+|C|))$. Moreover, whenever a valid covering exists, the algorithm outputs it by interpreting the corresponding encoding; otherwise, it correctly reports that no covering is possible. $\Box$
To summarize, given a covering guess size $m$ and ground set $\mathfrak{R}_{p}(S)$, we form $\mathfrak{R}_{p}(C)$, and check whether we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ with an exhaustive approach (Lemma 4.5). If we cannot find a covering, we return bad-guess; we cannot cover $\mathfrak{R}_{p}(S)$ using $[m]$. If we find a solution, use the covering given by $\mathfrak{R}_{p}(C)$ to cover $\mathfrak{R}_{p}(S)$ using $m^{\prime}]$, where $m^{\prime}=m+\lceil m/p\rceil$. This gives a $(1+1/p)$-approximation for covering $\mathfrak{R}_{p}(S)$. Applying the Proposition 4.4, we can get an approximating factor of $(1+2/p)(1+1/p)$, which can be as close as we want to $1$ by setting $p$ large enough. Using binary search adds a multiplicative factor of $\log\boldsymbol{OPT}$ to the runtime we used for the exhaustive search (see Lemma 4.5). Hence, the overall running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$. We can conclude the following:
Theorem 4.6.
There is a PTAS for the Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
Proof. Given an instance $S$ of Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+2/p)(1+1/p)\leq 1+\epsilon$. Construct $\mathfrak{R}_{p}(S)$ by rounding up elements in $S$. We use binary search over $m$ to find $\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))$, where $\boldsymbol{OPT}_{p}(S)$ is the minimum $m$ such that $S$ can be covered using $\mathfrak{R}_{p}([m])$. For each candidate set $[m]$:
  1. (i).
    Construct $\mathfrak{R}_{p}(C)$ where $C=[m]$ as per Definition 4.3.
  2. (ii).
    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.1. Given that $C^{\prime}=[m^{\prime}]$ covers $\mathfrak{R}_{p}(S)$, it covers $S$ as well. Using Proposition 4.4, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)(1+2/p)\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
Lemma 4.2 and Theorem 4.6 imply the following Corollary.
Corollary 4.7.
There exists a PTAS for Telephone Broadcasting in $k$-cycle graphs. In particular, for a $k$-cycle graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
4.2 PTAS for $k$-Path Graphs
In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-path graphs. Consider a $k$-path graph $G$ with endpoints $s$ and $t$.
In what follows, we assume that $s$ is the designated source, holding the message at time 0, and that $t$ acts as a late source, receiving the message at time $\alpha$, with $\alpha\leq d(s,t)$, where $d(s,t)$ denotes the distance between $s$ and $t$ in $G$. We show that any instance of the (single-source) problem with an arbitrary originator can be reduced to this double-source setting (with $s$ as the source and $t$ as the late source) without changing the approximation factor (Lemma 4.8).
In Lemma 4.10, we establish an equivalence between double-source broadcasting and the integer covering problem that we call Double Prefix Covering (Definition 4.9).
Finally, by applying a rounding-and-covering approach similar to the one used for Prefix Covering in $k$-cycle graphs, we obtain a PTAS for Double Prefix Covering, and consequently for Telephone Broadcasting in $k$-path graphs (Section 4.2.1). In what follows, we describe these steps in detail. First, we prove the following lemma, which implies we can restrict our attention to the double-source broadcasting.
Lemma 4.8.
Any $c$-approximation algorithm for Telephone Broadcasting in $k$-path graphs in the double-source setting—where $s$ is the primary source and $t$ is a late source receiving the message at time $\alpha$—can be extended to a $c$-approximation algorithm for the (single-source) general setting, where the originator $v_{o}$ may be any vertex of the graph.
Proof Sketch: Without loss of generality, assume $v_{o}$ is closer to $s$ than to $t$. At round $1$, $v_{o}$ sends the message along its path toward $s$, and at round $2$, it sends the message along the path toward $t$. Once $s$ receives the message, it immediately forwards it along its shortest path toward $t$. Finally, when $t$ receives the message, it first transmits it back along its path toward $v$ (to ensure that the neighbor is informed if it has not already been). These enforced decisions affect the broadcast time only by a constant number of rounds (the approximation remains the same). If $s$ is informed at time $\tau$, then $t$ is informed at time $\tau^{\prime}=\tau+\alpha$. $\Box$
Proof. Let $v_{o}$ be the originator vertex lying in path $P_{i}$ at distances $d(v_{o},s)$ and $d(v_{o},t)$ from endpoints $s$ and $t$, respectively. Without loss of generality, assume $d(v_{o},s)\leq d(v_{o},t)$. Let $d(v_{o},s)$. The message is first forwarded along the shortest path from $v_{o}$ to $s$, which takes $d(v_{o},s)$ rounds. At round $d(v_{o},s)+1$, vertex $s$ is informed and can begin broadcasting. Vertex $t$ is informed either directly through $v_{o}$ (after $d(v_{o},t)$ rounds) or through $s$ via some other path $P_{j}$ (whichever occurs first). Let $\alpha$ denote the delay with which $t$ receives the message relative to $s$. Once both $s$ and $t$ are informed, they proceed with the $c$-approximation algorithm on the remaining graph.
We consider two cases depending on how $t$ receives the message:
Case I: $t$ is informed directly through $v_{o}$. Here $\alpha=d(v_{o},t)-d(v_{o},s)+1$. If the global broadcast time is determined by a vertex in $P_{i}$, then the scheme is optimal. Otherwise, the process completes within $d(v_{o},s)+c\cdot\boldsymbol{br}^{-}(\alpha)$ rounds, where $\boldsymbol{br}^{-}(\alpha)$ is the optimal broadcast time for the subproblem excluding $P_{i}$ with sources $s$ and $t$ where $t$ has delay $\alpha$. We have $\boldsymbol{br}\geq d(v_{o},s)+\boldsymbol{br}^{-}(\alpha-1)\geq d(v_{o},s)-1+\boldsymbol{br}^{-}(\alpha)$ because $v_{o}$ can inform its neighbor toward $t$ in its first round, which results in $t$ getting informed one round sooner. Therefore, this yields the same approximation factor $c$.
Case II: $t$ is informed through $s$ via another path $P_{j}$. Let $\ell_{j}$ be the number of vertices of path $P_{j}$ (excluding $s$ and $t$). Then $t$ is informed at round $d(v_{o},s)+1+\ell_{j}$, giving delay $\alpha=\ell_{j}$. If the global broadcast time is determined by a vertex in $P_{i}$ or $P_{j}$, then the scheme is optimal. Otherwise, the process completes within $d(v_{o},s)+1+c\cdot\boldsymbol{br}^{-}(\alpha)$ rounds, where $\boldsymbol{br}^{-}(\alpha)$ is the optimal broadcast time for the subproblem excluding $P_{i}$ and $P_{j}$ with sources $s$ and $t$ where $t$ has delay $\alpha$. Since $\boldsymbol{br}\geq d(v_{o},s)+1+\boldsymbol{br}^{-}(\alpha)$, this yields the same approximation factor $c$. $\Box$
We introduce the following problem, which is the key part of our PTAS algorithms.
Definition 4.9.
The input to the Double Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers together with a non-negative integer $\beta$. The objective is to determine the minimum integer $m$ such that there exist functions $c:S\to\{0\}\cup[m]$ and $d:S\to\{0\}\cup[m-\beta]$ satisfying $c(i)+d(i)\geq\ell_{i}$ for all $i\in[n]$. Moreover, injectivity is required in the following sense: for any $i\neq j$, if both $c(i)>0$ (respectively $d(i)>0$), then $c(i)\neq c(j)$ (respectively $d(i)\neq d(j)$). In this setting, we say that each $\ell_{i}$ is covered by the pair $\{c(i),d(i)\}$, and we refer to $C=[m]$ and $D=[m-\beta]$ as the covering sets of $S$.
Example 8.
When $S=\{8,4,4,2\},$ the optimal solution is given by $C=[5]$ and $\beta=2$, and we have $c(1)=5,d(1)=3$; $c(2)=4,d(2)=0$, $c(3)=2,d(3)=2$, $c(4)=1,d(4)=1$.
In what follows, we establish the equivalence between double-source broadcasting in $k$-path graphs and Double Prefix Covering (Definition 4.9).
Lemma 4.10.
Let $G$ be a $k$-path graph formed by $k$ paths with common endpoints $s$ and $t$, where the $i$th path contains $\ell_{i}\geq 0$ internal vertices (in addition to $s$ and $t$). Consider a double-source broadcasting instance on $G$, where $s$ receives the message at time 0 and $t$ at time $\alpha$. Now, define an instance $I$ of Double Prefix Covering as $I=(S=\{\ell_{1},\ldots,\ell_{k}\},\;\beta=\alpha-1)$. Then computing the minimum broadcast time for this double-source Telephone Broadcasting instance is equivalent to solving the Double Prefix Covering instance $I$.
Proof. Consider the covering of the instance $I$ with parameter $m$. We claim that there exists a valid broadcasting scheme for $G$ within $\boldsymbol{br}=m$ rounds. For each $i\in[k]$, let $c(i)$ and $d(i)$ denote the assigned values in the covering. Then $s$ informs the $i$th path at round mc(i)+1m-c(i)+1 and $t$ at round md(i)+1m-d(i)+1. Thus, $s$ (respectively $t$) uses disjoint rounds drawn from $\{\,m-c+1\mid c\in C\,\}=[m]$ (respectively from $\{\,m-d+1\mid d\in[m-\alpha+1]\,\}=[\alpha,m]$). Hence, both $s$ and $t$ follow a valid broadcast scheme.
It remains to argue coverage. Since $c(i)+d(i)\geq\ell_{i}$, the remaining time after the calls at rounds mc(i)+1m-c(i)+1 and md(i)+1m-d(i)+1 is sufficient to cover all $\ell_{i}$ internal vertices of path $i$. Thus, the covering implies a feasible broadcasting scheme.
Conversely, suppose we have a broadcasting scheme using $\boldsymbol{br}$ rounds. For each $\ell_{i}\in S$, if $s$ and $t$ call vertices in path $i$ at rounds $y^{i}_{s}$ and $y^{i}_{t}$, define $c(i)=\boldsymbol{br}-y^{i}_{s}+1$ and $d(i)=\boldsymbol{br}-y^{i}_{t}+1$. Since calls at different paths use distinct rounds, the values $c(i)$ (resp. $d(i)$) are all distinct. Moreover, the chosen values lie in $\{\boldsymbol{br}-x+1\mid x\in[\boldsymbol{br}]\}=[\boldsymbol{br}]$ and $\{\boldsymbol{br}-x+1\mid x\in[\alpha,\boldsymbol{br}]\}=[\boldsymbol{br}-\beta]$, respectively. Finally, feasibility of the broadcasting scheme ensures $c(i)+d(i)=(\boldsymbol{br}-y^{i}_{s}+1)+(\boldsymbol{br}-y^{i}_{t}+1)\geq\ell_{i}$, so the functions $c$ and $d$ indeed form a valid covering of $S$. This establishes the equivalence. $\Box$
4.2.1 PTAS for Double Prefix Covering
In the light of Lemmas 4.8 and 4.10, presenting a PTAS for broadcasting in $k$-path graphs reduces to presenting a PTAS for the Double Prefix Covering problem. This is the focus of the present section. Specifically, we prove Theorem 4.13, which establishes the existence of a $(1+\epsilon)$-approximation algorithm for Double Prefix Covering for any $\epsilon>0$. Our approach closely parallels the rounding and covering techniques used in Section 4.1.1 for $k$-cycle graphs. In particular, it also addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the two covering sets into simplified versions with constant support (i.e., bounded numbers of distinct elements), so that exhaustive search becomes computationally feasible while still preserving provable approximation guarantees.
Our PTAS for Double Prefix Covering proceeds by reducing the original instance to one with bounded support. Given an input multiset $S$, we first form its rounded version $\mathfrak{R}_{p}(S)$, where each element is rounded up to a larger element of $S$ so that the number of distinct values becomes $p$. Here $p$ depends only on $\epsilon$. Although rounding inflates item sizes, we show that the optimal covering cost for $\mathfrak{R}_{p}(S)$ is at most a $(1+1/p)$ factor larger than that of $S$ (Proposition 4.11). Thus, solving the rounded instance suffices for approximation.
To search for the optimal covering, we use binary search over candidate values $m$, where $m$ determines the covering sets $C=[m]$ and $D=[m-\beta]$. For each candidate $m$, we construct rounded sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, each with support size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, then the original instance can be covered using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ with $m^{\prime}\leq(1+1/p)m$ (4.2). Conversely, if we cannot cover $\mathfrak{R}_{p}(S)$ with the rounded sets, then it is not possible to cover it using $[m]$ and $m-\beta]$. It therefore suffices to test whether $\mathfrak{R}_{p}(S)$ is coverable using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$. Since all three sets have constant support (bounded by $p$), this check can be carried out by exhaustive search in constant time (Lemma 4.12). The remainder of this section provides details of the algorithm.
Proposition 4.11.
$\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq\boldsymbol{OPT}(S)+\lceil|S|/p\rceil$, where $\boldsymbol{OPT}$ is the optimal cover for a specific instance of Double Prefix Covering.
Proof. By definition, $\mathfrak{R}_{p}(S)$ consists of $p$ parts. Given an optimal cover of $S$, form a cover of $\mathfrak{R}_{p}(S)$ by assigning the numbers given to the part $i$ in $S$ to the part $(i+1)$ in $\mathfrak{R}_{p}(S)$ for $i\in[p-1]$. As the size and value of part $i+1$ is not greater than the value of part $i$, the set of numbers that can cover part $i$ can also cover part $i+1$.
Therefore, the first part is the only part that needs to be covered. We show that adding an extra $\lceil|S|/p\rceil$ to the covering set is enough to cover the first part. This is because every $\ell_{i}\in S$ is less than $2\boldsymbol{OPT}(S)-\beta$, as otherwise it would not be possible to cover $v_{i}$ with a number in $C$ and a number in $D$ for the optimal value of $m$. Thus, by adding an extra $\lceil|S|/p\rceil$ items to the $m$, we can cover $\mathfrak{R}_{p}(S)$. $\Box$
Since $\boldsymbol{OPT}(S)\geq|S|$, it follows that any covering of $\mathfrak{R}_{p}(S)$ yields a $(1+1/p)$-approximation for covering $S$. This allows us to restrict attention to $\mathfrak{R}_{p}(S)$, which has constant support $p$.
To cover $\mathfrak{R}_{p}(S)$, we apply a binary strategy. For a candidate value $m$, we design a procedure that either produces a covering using $m+\lceil m/p\rceil]$, or certifies that no covering exists with $[m]$. By iteratively refining $m$, we identify the smallest value for which $\mathfrak{R}_{p}(S)$ can be covered using the rounded covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, where $C=[m]$ and $D=[m-\beta]$.
Claim 4.2.
Given that we can cover $S$ with $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, we can cover $S$ using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$, where $m^{\prime}=m+\lceil m/p\rceil$.
Proof. Based on definition, $\mathfrak{R}_{p}(C)$ contains $p$ parts. Suppose that we have a covering with $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ for $S$. Associate value $m+\lceil m/p\rceil-(i-1)$ to the $i$’th largest number in $C$ for $i\in[m]$. This transformation only increases the value as the first block with original value $m$ matches to $m+1,\ldots,m+\lceil m/p\rceil$, the second block matches to $m-\lceil m/p\rceil+1,\ldots,m$, and so on. More specifically, the $i$’th block matches to $m-(i-1)\lceil m/p\rceil+1,\ldots,m-(i-2)\lceil m/p\rceil$, which is greater than the original value of the block. Similarly, associate value $m+\lceil m/p\rceil-\beta-(i-1)$ to the $i$’th largest number in $D$ for $i\in[m-\beta]$. As the new sets are subsets of $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$, and every value in $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ is now larger than its original value, the same covering can cover $S$ with the corresponding values in $C^{\prime}$ and $D^{\prime}$. $\Box$
Lemma 4.12.
Given three sets $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|+|D|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ or correctly reports that no such covering is possible.
Proof. The ground set $\mathfrak{R}_{p}(S)$ and the covering sets have constant support $p$. For any fixed distinct member of $\mathfrak{R}_{p}(S)$ like $v$ and each combination pair of members of $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, we can encode the number of times the pair is used to cover the members of $\mathfrak{R}_{p}(S)$ with value $v$ in $\log|S|$ bits. This is because we have at most $|S|$ members with value equal to $v$.
Given that there are $p^{2}$ distinct pairs from the covering set and $p$ types of value in $\mathfrak{R}_{p}(S)$, any potential solution for the problem can be encoded in $p^{2}.p.\log|S|=p^{3}\log|S|$ bits; we can exhaustively try all and answer whether $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ by trying all potential solutions.
Thus, the algorithm tries $\mathcal{O}(2^{p^{3}}|S|)$ solutions. And, each solution takes linear time to check. Therefore, overall it takes $(2^{p^{3}}|S|(|S|+|C|+|D|))$ to decide whether it is possible to cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$. $\Box$
To summarize, for a candidate covering size $m$ and ground set $\mathfrak{R}_{p}(S)$, we construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ and test, via exhaustive search (Lemma 4.12), whether $\mathfrak{R}_{p}(S)$ can be covered. If no covering is found, we conclude that $\mathfrak{R}_{p}(S)$ cannot be covered with $[m]$ (bad guess). If a covering is found, we extend it to a covering with $m^{\prime}=m+\lceil m/p\rceil$, thereby obtaining a $(1+1/p)$-approximation for $\mathfrak{R}_{p}(S)$. Combining this with Proposition 4.11, we achieve an overall approximation factor of $(1+1/p)^{2}$, which approaches 1 as $p$ grows.
Theorem 4.13.
There is a PTAS for the Double Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
Proof. Given an instance $S$ of Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+1/p)^{2}\leq 1+\epsilon$. Construct $\mathfrak{R}_{p}(S)$ by rounding up elements in $S$. We use binary search over $m$ to find $\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))$, where $\boldsymbol{OPT}_{p}(S)$ is the minimum $m$ such that $S$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. For each candidate set $[m]$:
  1. (i).
    Construct $\mathfrak{R}_{p}(C)$ where $C=[m]$ as per Definition 4.3.
  2. (ii).
    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.2. Given that $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ cover $\mathfrak{R}_{p}(S)$, they cover $S$ as well. Using Proposition 4.11, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)^{2}\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
Combining Lemma 4.10 and Theorem 4.13 implies the following Corollary.
Corollary 4.14.
There exists a PTAS for Telephone Broadcasting in $k$-path graphs. In particular, for a $k$-path graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
5 Graphs of Bounded Bandwidth
In this section, we present a polynomial algorithm that computes the optimal broadcasting time for graphs of bounded bandwidth. Intuitively, the bandwidth constraint limits the degree of $G$ (in contrast to other families such as $k$-cycle or $k$-path graphs), and a bandwidth ordering exposes a small “active window” of vertices at each step. This property enables a dynamic programming (DP) approach. For the remainder of this section, we assume that a bandwidth-$k$ ordering $(v_{1},\ldots,v_{n})$ of the vertices is given. For clarity of exposition, we also assume that the source of the broadcast is $v_{1}$; this restriction will be relaxed later.
DP Table. We define a DP table with entries of the form $dp[i][s]$, where $i\in[0,n]$ and $s$ is a state. The value $dp[i][s]$ denotes the minimum broadcast time for informing the first $i$ vertices in the ordering while ending in configuration $s$. The state $s$ encodes all necessary information about the boundary vertices $B_{i}=\{v_{i-k+1},\ldots,v_{i}\}$, namely, how these vertices are informed and how they connect to the rest of a potential broadcast tree. Intuitively, a state describes the partial structure of a broadcast forest spanning boundary vertices together with their timing information.
We fill the table row by row, starting from $i=k$ up to $i=n$. The last row corresponds to the entire graph; the minimum value among its states equals the optimal broadcast time.
Definition of states. Fix $i$, and let $u_{1},\ldots,u_{k}$ denote the boundary vertices. Let $Bf=\{x_{1},\ldots,x_{k}\}$ be the $k$ vertices immediately before $B_{i}$ and $Af=\{y_{1},\ldots,y_{k}\}$ be the $k$ vertices immediately after $B_{i}$. Together, the vertices $x_{1},\ldots,x_{k},\;u_{1},\ldots,u_{k},\;y_{1},\ldots,y_{k}$ form the active window. Every neighbor of a boundary vertex lies within this window.
A state $s$ encodes a potential partial broadcast tree consistent with the active window by storing the following information:
  • Parents. For each $u_{i}$, its parent $p(u_{i})$ (unless it is the source $v_{1}$). Each $p(u_{i})$ is chosen from at most $2k$ active vertices or a special “no parent” symbol, yielding $(2k+1)^{k}$ possibilities.
  • Ranks. For each $u_{i}$, its rank among the children of $p(u_{i})$, ordered by the rounds in which $p(u_{i})$ calls them. Since $p(u_{i})$ has at most $2k$ children, there are at most $2k$ options per vertex, and $(2k)^{k}$ total options.
  • Outgoing calls. For each $u_{i}$, the active vertex that it calls in each round after being informed. For every round slot, there are $2k$ options (one of up to 2k12k-1 children or a null option if unused), giving at most $(2k)^{2k+1}$ possibilities per $u_{i}$.
  • Inform times. The time at which the root of each subtree is informed. Since there are at most $k$ subtrees (each containing a boundary vertex), and each root may be at a round in $[n]$, this yields $n^{k}$ possibilities.
Proposition 5.1.
If $S$ denotes the space of all states, then $|S|\;\leq\;(2k+1)^{k}\cdot(2k)^{k}\cdot(2k)^{k(2k+1)}\cdot n^{k}$, which is $\mathcal{O}(n^{k})$ for constant $k$.
For a state $s$, let $\ell(s)$ denote the latest time at which any active node in the broadcast subtrees of $s$ is informed. Note that the global broadcast time of a scheme realizing $s$ may exceed $\ell(s)$ if the last informed vertex lies outside the active window.
Filling the DP table. A state $s$ is valid for step $i$ if its parent–child assignments among nodes in its broadcast forest respect the edges of $G$. Also, the outgoing calls in the state must match the rank and parent fields. For example, if $u_{1}$ and $u_{2}$ are not adjacent in $G$ but $u_{1}$ is the parent of $u_{2}$ in $s$, then $s$ is invalid at $i$. For any invalid state at step $i$, we set $dp[i][s]=+\infty$.
Base case. When $i=k$, any valid state represents a single broadcast tree rooted at $v_{1}$ (informed at time 0). For such states, we set $dp[k][s]=\ell(s)$.
Transitions. To compute $dp[i+1][s]$ for a valid state $s$, we look for states $s^{\prime}$ that are valid at step $i$ and compatible with $s$. Compatibility means that information concerning the overlap between active states at steps $i$ and $i+1$ matches in both states: the inform times of shared vertices, their parent–child relations, and the directions and timings of edges incident to them must all agree. Further, indices shift in the natural way: for example, $u_{1}$ at step $i$ corresponds to $x_{k}$ at step $i+1$, and $u_{2}$ at step $i$ corresponds to $u_{1}$ at step $i+1$. Vertices $x_{1}$ and $y_{k}$ enter or leave the active window accordingly. See Figure 4 for an illustration.
If $s^{*}$ is a compatible state with $s$ with minimum $dp[i][s^{*}]$, we set $dp[i+1][s]\;=\;\max\{dp[i][s^{*}],\ell(s)\}.$ We also record $\mathrm{previous}(s)=s^{*}$, which allows reconstruction of the optimal broadcast tree. If no compatible state exists, then $dp[i+1][s]=+\infty$. Intuitively, realizing state $s$ at $i+1$ requires extending from some compatible predecessor state at $i$, and the broadcast time is determined by whichever is larger: the time already incurred or the latest inform time among the active vertices.
Example 9.
In Figure 4, suppose that at step $i+1$, the best compatible predecessor of state $S_{2}$ is $S_{1}$ with $dp[i][S_{1}]=120$. Since $\ell(S_{2})=109$, we obtain $dp[i+1][S_{2}]=\max\{120,109\}$.
[…]
Figure 4: An illustration of states and their compatibility for a graph of bandwidth $k=3$. The bandwidth ordering is $(\ldots,x_{1},x_{2},x_{3},u_{1},u_{2},u_{3},y_{1},y_{2},y_{3},\ldots)$. The boundary vertices $u_{i}$ are highlighted in red. When computing DP values at step $i+1$, state $S_{2}$ is compatible with state $S_{1}$. Note that the shift in the active window implies, for example, that $u_{1}$ in $S_{1}$ at step $i$ becomes $x_{3}$ in $S_{2}$ at step $i+1$, $u_{2}$ becomes $u_{1}$, $x_{2}$ becomes $x_{1}$, and so on.
Sources other than $v_{1}$. If the broadcast source is not $v_{1}$, then we consider all $O((2k)!)$ possible neighbor orderings in which the source $s$ informs its neighbors. For a fixed ordering, the problem splits into two independent subproblems: one for vertices preceding $s$ in the bandwidth ordering and one for vertices following it. The DP applies symmetrically to both sides. The only modification is in the base case: a state is valid only if it is consistent with the fixed ordering of $s$’s initial calls. Since $k$ is constant, this does not affect the asymptotic time complexity.
Using an inductive argument, one can establish the correctness of the above approach.
Theorem 5.2.
Telephone Broadcasting can be solved optimally on graphs of bandwidth $k$ in time $f(k)\cdot n^{k^{2}+O(k)}$. In particular, for a constant $k$, there is a polynomial-time algorithm for computing an optimal broadcast schedule.
Proof.
Setup. Recall that we have a fixed bandwidth-$k$ ordering $(v_{1},\dots,v_{n})$. As discussed, at step $i$ the DP keeps states $s$ that summarize the broadcast restricted to the active window $W_{i}:=\{x_{1},\dots,x_{k},u_{1},\dots,u_{k},y_{1},\dots,y_{k}\}$, where $B_{i}=\{u_{1},\dots,u_{k}\}$ is the set of boundary nodes. Recall that, for a state $s$, $\ell(s)$ is the latest inform time among active vertices encoded by $s$.
Correctness. We use induction to prove the following invariant:
Inductive invariant. For every $i\geq k$ and every valid state $s$ at step $i$, the entry $dp[i][s]$ of the DP table is the minimum broadcast time of any feasible broadcast scheme for informing $v_{1},\ldots,v_{i}$ which ends up at state $s$.
Base $i=k$. In the base case, the round in which each vertex of $v_{1},\ldots,v_{k}$ is informed is embedded into the state $s$. This is because they are all boundary vertices. Therefore, the minimum broadcast time over schedules realizing $s$ is exactly $\ell(s)$, so $dp[k][s]=\ell(s)$ as set by the algorithm.
Step $i\to i+1$. Let $s$ be any valid state at step $i+1$. Choose a compatible predecessor $s^{*}$ at step $i$ that minimizes $dp[i][s^{*}]$.
By the inductive hypothesis there is broadcast scheme $\mathcal{S}^{*}$ that realizes $s^{*}$ and informs $v_{1},\ldots,v_{i}$ in $dp[i][s^{*}]$ rounds. Extend $\mathcal{S}^{*}$ inside the active window to realize $s$ at step i+1i{+}1 (this is possible by compatibility and validity). The broadcast time of the extended scheme is $\max(dp[i][s^{*}],\ell(s))$, which is exactly the value assigned by the recurrence. Hence, dp[i+1][s]dp[i{+}1][s] upper-bounds the true optimum among schedules realizing $s$.
On the other hand, let $\mathcal{S}$ be any feasible schedule on $G_{i+1}$ that realizes $s$ with minimum broadcast time $\tau$. Let $s^{\prime}$ be the state of $\mathcal{S}$ at step $i$; then $s^{\prime}$ is valid at $i$ and compatible with $s$. By the inductive hypothesis, $dp[i][s^{\prime}]$ is no more than the broadcast time of $\mathcal{S}$ for informing $v_{1},\ldots,v_{i}$. Since realizing $s$ from $s^{\prime}$ cannot finish later than $\max\{dp[i][s^{\prime}],\ell(s)\}$, we get $dp[i{+}1][s]\leq\max\{dp[i][s^{\prime}],\ell(s)\}\leq\tau.$ Thus, dp[i+1][s]dp[i{+}1][s] lower-bounds the optimum as well. This proves the inductive invariant.
We conclude that taking the minimum $dp[n][s]$ over valid states $s$ yields the optimal broadcast time on $G$.
Sources other than $v_{1}$. If the source is $s\neq v_{1}$, we enumerate the $O((2k)!)$ possible orders of $s$’s calls (only neighbors within distance $k$ in the bandwidth ordering matter). For each fixed order, the same inductive argument applies independently to the two sides of $s$, and we take the best over the constant many orders. This preserves optimality.
Running time. Let $S$ be the set of all states. By Proposition 5.1, we have $|S|\leq(2k{+}1)^{k}\cdot(2k)^{k}\cdot(2k)^{k(2k+1)}\cdot n^{k}\;=\;g(k)\,n^{k}.$ That is, the size of the DP table would be at most $g(k)n^{k+1}$. For every $i$ and $s$, comparing every state $s^{\prime}$ of the previous iteration (i.e. i1i-1) to find out whether they are compatible would take $O(|S|.n)$ time. Therefore, the cost of computing dp for all states is $(g(k))^{2}\,n^{2k+2}$. Overall, this yields a total time $f(k)\cdot n^{O(k)}$, for some function $f(k)$ independent of $n$. Note that the extra $O((2k)!)$ factor for non-$v_{1}$ sources is absorbed into $f(k)$ since $k$ is constant. This matches the claimed bound. $\Box$
6 Concluding Remarks
In this work, we advanced the study of Telephone Broadcasting by establishing new hardness results and developing approximation schemes for specific graph families. On the hardness side, we resolved long-standing questions by proving NP-hardness for $k$-cycle and $k$-path graphs. On the algorithmic side, we presented polynomial-time approximation schemes (PTAS) for Telephone Broadcasting in both $k$-cycle and $k$-path graphs. We also introduced an exact algorithm for graphs of bounded bandwidth, thereby extending the tractable frontier of the problem.
As future work, it would be interesting to adapt the integer covering formulations to other broadcasting models or broader sparse graph classes. In particular, for series–parallel graphs, our hardness result via $k$-path graphs rules out tractability, yet, to the best of our knowledge, no constant-factor approximation is currently known.

### v2

@@
-On Hardness and Approximation of Broadcasting in Sparse Graphs
+On Hardness and Approximation of Broadcasting in Structured Graphs
 Abstract
-We study the Telephone Broadcasting problem in sparse graphs.
-Given a designated source in an undirected graph, the task is to disseminate a message to all vertices in the minimum number of rounds, where in each round every informed vertex may inform at most one uninformed neighbor.
+We study the Telephone Broadcasting problem in graphs with restricted structure.
+Given a designated source in an undirected graph, the goal is to disseminate a message to all vertices in the minimum number of rounds, where in each round every informed vertex may inform at most one uninformed neighbor.
 For general graphs with $n$ vertices, the problem is NP-hard.
-Recent work shows that the problem remains NP-hard even on restricted graph classes such as cactus graphs of pathwidth $2$ [Aminian et al., ICALP 2025] and graphs at distance-1 to a path forest [Egami et al., MFCS 2025].
+Recent work shows that the problem remains NP-hard even on restricted graph classes such as graphs of treewidth 2 [Tale 2025], cactus graphs of pathwidth $2$ [Aminian et al., ICALP 2025] and graphs at distance-1 to a path forest [Egami et al., MFCS 2025].
 In this work, we investigate the problem in several sparse graph families.
-We first prove NP-hardness for $k$-cycle graphs, namely graphs formed by $k$ cycles sharing a single vertex, as well as $k$-path graphs, namely graphs formed by $k$ paths with shared endpoints.
-Despite multiple efforts to understand the problem in these simple graph families, the computational complexity of the problem had remained unsettled, and our hardness results answer open questions by Bhabak and Harutyunyan [CALDAM 2015] and Harutyunyan and Hovhannisyan [COCAO 2023] concerning the problem’s complexity in $k$-cycle and $k$-path graphs, respectively.
-On the positive side, we present Polynomial-Time Approximation Schemes (PTASs) for $k$-cycle and $k$-path graphs, improving over the best existing approximation factors of $2$ for $k$-cycle graphs and an approximation factor of $4$ for $k$-path graphs.
-Moreover, we identify a structural frontier for tractability by showing that the problem is solvable in polynomial time on graphs of bounded bandwidth.
-This result generalizes existing tractability results for special sparse families such as necklace graphs.
+We first prove NP-hardness for cycle-star graphs, namely graphs formed by $k$ cycles sharing a single vertex, as well as melon graphs, namely graphs formed by $k$ paths with shared endpoints.
+Despite multiple efforts to understand the problem in these simple graph families, the computational complexity of the problem remained unsettled.
+Our hardness results answer open questions by Bhabak and Harutyunyan [CALDAM 2015] and Harutyunyan and Hovhannisyan [COCOA 2023] concerning the problem’s complexity in cycle-star and melon graphs, respectively.
+On the positive side, we present Efficient Polynomial-Time Approximation Schemes (EPTASs) for cycle-star and melon graphs, improving over the best existing approximation factors of $2$ for both graph families.
+Moreover, we identify a structural frontier for tractability by showing that the problem is solvable in polynomial time on graphs of bounded cutwidth, a class that generalizes other families such as graphs of bounded bandwidth.
+This result subsumes existing tractability results for special sparse graph families such as necklace graphs.
+Finally, for split graphs, a fundamental class of highly structured graphs, we obtain a polynomial-time algorithm with approximation factor 1.76.
+This improves on the previously known factor $2$ bound; the same approach also applies to the multi-source setting.
 1 Introduction
@@
 On the other hand, Elkin and Kortsarz [10] designed an $\mathcal{O}(\log n/\log\log n)$-approximation algorithm for general graphs, which is the best existing approximation algorithm.
-Whether a constant-factor approximation exists for general graphs remains open, though such results are known for specific classes, including unit disk graphs [18] and cactus graphs [1].
+Whether a constant-factor approximation exists for general graphs remains open, though such results are known for specific classes, including unit disk graphs [24], cactus graphs [2], and graphs of bounded pathwidth [2, 8].
 A substantial body of work has investigated the complexity of Telephone Broadcasting on restricted graph classes.
-On the positive side, polynomial-time algorithms are known for several families, including trees [10], grids, stars of cliques, fully connected trees, chains of rings, necklace graphs, and certain subclasses of cactus graphs; see, e.g., [5, 12, 17, 20, 9].
-On the negative side, hardness results persist even for structurally simple graphs: the problem is NP-hard in planar cubic graphs and in graphs of bounded treewidth.
-More recently, Tale showed NP-hardness for graphs of pathwidth $3$, while Aminian et al.
-[1] established hardness for so-called snowflake graphs (cactus graphs of pathwidth $2$), resolving a long-standing open question about the complexity of broadcasting in cacti.
+On the positive side, polynomial-time algorithms are known for several families, including trees [12], grids, stars of cliques, fully connected trees, necklace graphs (chains of rings), and certain subclasses of cactus graphs; see, e.g., [7, 14, 23, 26, 11].
+On the negative side, hardness results persist even for structurally simple graphs such as planar cubic graphs [19].
+More recently, Tale [27] showed that the problem remains NP-hard on graphs at distance 2 from path forests, as well as graphs of pathwidth 3 and treewidth 2.
+Aminian et al.
+[2] established hardness for so-called snowflake graphs (cactus graphs of pathwidth $2$).
 Even stronger results were obtained by Egami et al.
 [8], who proved NP-hardness for graphs at distance 1 from path forests, i.e., graphs in which deleting a single vertex yields a path forest.
-See Figure 1 for illustrations of snowflake graphs and graphs at distance 1 from path forests.
-These results highlight intriguing frontier cases where the complexity of the problem remains unresolved.
-One such case is that of $k$-cycle graphs (also called flower graphs), obtained by joining $k$ cycles at a single vertex.
+Figure 1 illustrates several hard sparse graph instances.
+These results highlight intriguing frontier cases where the problem’s complexity remains unresolved.
+One such case is that of cycle-star graphs (also called flower graphs), obtained by joining $k$ cycles at a single vertex.
 Despite their simple definition, these graphs exhibit surprisingly rich behavior.
-Broadcasting on $k$-cycle graphs has been studied from an approximation perspective: the best known algorithm achieves a factor of 2 [2], yet the exact complexity remains open.
-A closely related family is that of $k$-path graphs, formed by joining $k$ paths at their endpoints.
-Several approximation algorithms are known for $k$-path graphs [3, 14], with the current best result being a 2-approximation [14].
+Telephone Broadcasting on cycle-star graphs has been studied from an approximation perspective: the best known algorithm achieves a factor of 2 [3], yet the exact complexity remains open.
+A closely related family is that of melon graphs, formed by joining $k$ paths at their endpoints.
+Several approximation algorithms are known for melon graphs [4, 15], with the current best result being a 2-approximation [15].
 Nevertheless, the precise computational complexity of the problem for this family also remains open.
+Another structurally interesting graph family is that of split graphs, which consist of a clique and an independent set and therefore fall outside the sparse regime.
+For single-source Telephone Broadcasting, a 2-approximation algorithm is known for split graphs [16].
+Moreover, the multi-source variant of Telephone Broadcasting is NP-hard even on split graphs [20], underscoring that hardness persists despite strong structural restrictions.
 1.1 Contribution
 Our contributions in this paper are as follows. In a nutshell, we tighten the frontier between the tractable and intractable cases of the Telephone Broadcasting problem.
-  • We prove that Telephone Broadcasting is NP-Complete in two of the simplest sparse graph families: $k$-cycle graphs (Theorem 3.4) and $k$-path graphs (Theorem 3.5).
+  • We prove that Telephone Broadcasting is NP-Complete in two of the simplest sparse graph families: cycle-star graphs (Theorem 3.4) and melon graphs (Theorem 3.5).
 Both families have been studied in the broadcasting literature, yet their complexity status remained open.
 Our hardness results also imply hardness for broader classes, including series–parallel graphs (SP-graphs), for which no hardness was previously known.
-  • We design polynomial-time approximation schemes (PTAS) for both $k$-cycle graphs (Corollary 4.7) and $k$-path graphs (Corollary 4.14).
-To the best of our knowledge, these are the first graph families for which Telephone Broadcasting admits both a hardness result and a matching PTAS, highlighting their role as boundary cases between tractability and intractability.
-  • We present a polynomial-time algorithm for optimal broadcasting in graphs of bounded bandwidth (Theorem 5.2).
-This result generalizes several known tractable cases, such as chains of rings and related structures.
-Notably, Telephone Broadcasting is among the few problems exhibiting a complexity gap between graphs of bounded bandwidth (where the problem is tractable) and graphs of bounded pathwidth (where the problem is NP-hard).
-In particular, $k$-cycle and $k$-path graphs both have pathwidth 2 but unbounded bandwidth, placing them precisely on the intractable side of this gap.
-(a) A snowflake graph (cactus of pathwdith 2) [1]
-(b) Graph of distance 1 to path forest [6]
-(c) A $k$-cycle graph (Theorem 3.4)
-(d) A $k$-path graph (Theorem 3.5)
+  • We design efficient polynomial-time approximation schemes (EPTASs) for both cycle-star graphs (Corollary 4.7) and melon graphs (Corollary 4.14).
+To the best of our knowledge, these are the first known graph families for which Telephone Broadcasting admits both a hardness result and a matching EPTAS, highlighting their role as boundary cases between tractability and intractability.
+Our EPTASs extend naturally to the multi-source setting.
+  • We present a polynomial-time algorithm for optimal broadcasting in graphs of bounded cutwidth (Theorem 5.2).
+This result generalizes several previously known tractable cases, such as necklace graphs (chains of rings) and related structures [17], and also applies to the multi-source setting.
+Since graphs of bounded bandwidth also have bounded cutwidth,  [6], our result further implies tractability for graphs of bounded bandwidth.
+    Notably, Telephone Broadcasting is among the few problems exhibiting a complexity gap between graphs of bounded cutwidth (where the problem is tractable) and graphs of bounded pathwidth (where the problem is NP-hard [27]).
+In particular, cycle-star and melon graphs both have pathwidth 2 but unbounded cutwidth, placing them precisely on the intractable side of this gap.
+  • We present a $1.76$-approximation algorithm for Telephone Broadcasting on split graphs (Theorem 6.2), improving over the known factor 2 bound [16].
+Together with our tractability results for bounded-cutwidth graphs and hardness results for sparse families, this places split graphs as an intermediate structural class.
+The algorithm applies to both single-source and multi-source settings, the latter known as being APX-hard [20].
+(a) A graph of distance 2 to path forest [27]
+(b) A snowflake graph (cactus of pathwidth 2) [2]
+(c) Graph of distance 1 to path forest [8]
+(d) A cycle-star graph (Theorem 3.4)
+(e) A melon graph (Theorem 3.5)
 Figure 1: Sparse graph families for which Telephone Broadcasting is proven to be NP-hard.
 1.2 Overview of Techniques
-Our results combine hardness reductions, approximation schemes, and dynamic programming, each tailored to the structural properties of the graph families under consideration.
+Our results combine hardness reductions, approximation schemes, and dynamic programming, together with fixed-dimension integer programming and min-cost flow formulations, each tailored to the structural properties of the graph families under consideration.
 We give a brief outline here.
-  • For the negative results, we design reductions that encode scheduling constraints directly into the graph structure.
-In particular, the NP-hardness proofs for $k$-cycle and $k$-path graphs exploit the interaction between cycle (or path) lengths and the timing of broadcast calls.
-The shared vertex in $k$-cycles (or the pair of endpoints in $k$-paths) acts as the originator, and we show that deciding feasibility becomes equivalent to a combinatorial partitioning problem whose complexity propagates to the broadcasting instance.
-Our reductions build a more restricted variant of Numerical 3-Dimensional Matching (RN3DM) (Definition 3.1), previously employed in hardness proofs for telephone broadcasting.
-This restricted variant, which is known to be still NP-hard [22], makes it possible to encode number-theoretic constraints using only these simple graph families.
-  • We design polynomial-time approximation schemes for $k$-cycle and $k$-path graphs by reducing the broadcasting problem to covering problems over integers.
-For $k$-cycle graphs, the task reduces to covering a multiset of cycle lengths $\{\ell_{1},\ldots,\ell_{k}\}$ with the smallest possible broadcast round $\beta$ using pairs of numbers from $\beta\,]$, so that for each $\ell_{i}$ there exist $a,b\in[\,\beta\,]$ associated with $\ell_{i}$ such that $a+b\geq\ell_{i}$ (See Definition 4.1).
-For $k$-path graphs, the covering problem is slightly different: we seek the smallest $\beta$ such that each $\ell_{i}$ can be covered by one number from $\beta\,]$ and one from $\beta-d\,]$, where $d$ is the length of the shortest path (see Definition 4.9).
-In both cases, approximate solutions to the number-theoretic covering problems translate into near-optimal broadcasting schedules, yielding PTASs for $k$-cycle and $k$-path graphs.
-  • For finding optimal broadcast schemes in graphs of bounded bandwidth, we exploit the fact that a bandwidth-$k$ ordering restricts each vertex to neighbors within a sliding window of $\mathcal{O}(k)$ vertices.
-This allows us to describe the status of any partial broadcast tree using a compact boundary state and use a dynamic programming approach which has a table with $n$ rows (one for each prefix of the ordering) and at most $n^{\mathcal{O}(k^{2})}$ states per row, giving a total size bounded by $f(k)\cdot n^{\mathcal{O}(k^{2})}$.
+  • For the negative results, we design reductions that encode scheduling constraints directly into the graph structure, adapting and extending hardness techniques introduced by Tale [27] and Egami et al.
+[8].
+In particular, the NP-hardness proofs for cycle-star and melon graphs exploit the interaction between cycle (or path) lengths and the timing of broadcast calls.
+The shared vertex in cycle-star graphs (or the pair of endpoints in melon graphs) acts as the originator, and we show that deciding feasibility becomes equivalent to a combinatorial partitioning problem whose complexity propagates to the Telephone Broadcasting instance.
+Our reductions build a more restricted variant of Numerical 3-Dimensional Matching (RN3DM) (Definition 3.1), previously employed in hardness proofs for Telephone Broadcasting by Tale [27] and Egami et al.
+[8].
+The restricted variant that we use, which is known to be still NP-hard [29], makes it possible to encode number-theoretic constraints using only these simple graph families.
+  • We design efficient polynomial-time approximation schemes for cycle-star and melon graphs by reducing the Telephone Broadcasting problem to covering problems over integers.
+For cycle-star graphs, the task reduces to covering a multiset of cycle lengths $\{\ell_{1},\ldots,\ell_{k}\}$ with the smallest possible broadcast round $\beta$ using pairs of numbers from $\beta\,]$, such that for each $\ell_{i}$ there exist $a,b\in[\,\beta\,]$ with $a+b\geq\ell_{i}$ (see Definition 4.1).
+For melon graphs, the covering problem is slightly different: we seek the smallest $\beta$ such that each $\ell_{i}$ can be covered by one number from $\beta\,]$ and one from $\beta-d\,]$, where $d$ is the length of the shortest path between two endpoints (see Definition 4.9).
+In both cases, approximate solutions to the resulting number-theoretic covering problems translate into near-optimal broadcasting schedules; moreover, the rounded covering instances have constant support and can be solved exactly using fixed-dimension integer linear programming, yielding EPTASs for cycle-star and melon graphs.
+  • For finding optimal broadcast schemes in graphs of bounded cutwidth, we exploit the fact that a cutwidth-$k$ ordering restricts each vertex to neighbors within a sliding window of $\mathcal{O}(k)$ vertices.
+This allows us to describe the status of any partial broadcast tree using a compact boundary state and use a dynamic programming approach which has a table with $n$ rows (one for each prefix of the ordering) and at most $f(k)\cdot n^{k}$ states per row, giving a total size bounded by $f(k)^{2}\cdot n^{2k+1}$, in which $f(.)$ is a function of $k$ that is independent of $n$.
+  • For split graphs, we establish that informing independent-set vertices early is never beneficial (Lemma 6.1); equivalently, one can assume without loss of optimality that broadcasts to the independent set are deferred as long as possible.
+This structural insight allows us to decouple the broadcast process at each vertex into an initial phase within the clique, followed by a controlled dissemination to the independent set.
+As a result, the problem becomes related to constructing minimum-degree and minimum-overload coverings of the independent set by clique vertices.
+When discussing these covering problems, at a high level, we seek to minimize the number of clique vertices that are responsible for covering many independent-set vertices.
+Such high-load clique vertices may constitute the main bottleneck of the broadcast process and are therefore informed earlier, so that they can switch to informing independent-set vertices sooner than others.
+We compute a suitable covering via a reduction to a minimum-cost flow problem, yielding an approximation algorithm with approximation factor $1.76$ for both the single-source and multi-source settings.
 2 Preliminaries
-We use notation $[n]$ and $[n,m]$ for positive integers $n$ and $m$ for the set of all natural numbers less than or equal to $n$ and all natural numbers between $n$ and $m$ (inclusive), respectively (i.e.
+We use notation $[n]$ and $[n,m]$ for positive integers $n$ and $m$ for the set of all positive integers less than or equal to $n$ and all positive integers between $n$ and $m$ (inclusive), respectively (i.e.
 $n]=\{1,2,\ldots,n\}$ and $n,m]=\{n,n+1,\ldots,m\}$).
@@
 When the context is clear, we simply write $\boldsymbol{br}^{*}$} in place of ${\boldsymbol{br}^{*}}(G,s)$.
-Definition 2.2 ($k$-cycle and $k$-path graphs).
-A $k$-cycle graph $G$ contains $k$ cycles with exactly one common vertex $r$. A $k$-path graph $G$ contains $k$ paths with shared endpoints $s$ and $t$.
-See Figure 1 (parts (c) and (d)) for illustrations of $k$-cycle and $k$-path graphs.
-Observe that any $k$-cycle (respectively, $k$-path) graph can be fully described by a set of integers specifying the number of vertices on each cycle (respectively, path).
+We also consider a multi-source variant of Telephone Broadcasting, in which a set $S\subseteq V$ of vertices is initially informed at round 0.
+The broadcasting process proceeds as before, with each informed vertex informing at most one uninformed neighbor per round.
+The goal is again to minimize the number of rounds until all vertices are informed.
+Throughout the paper, unless explicitly stated otherwise, Telephone Broadcasting denotes the single-source broadcasting problem.
+Definition 2.2 (Cycle-star and Melon Graphs).
+A cycle-star graph $G$ contains $k$ cycles with exactly one common vertex $r$. A melon graph $G$ contains $k$ paths with shared endpoints $s$ and $t$.
+See Figure 1 (parts (c) and (d)) for illustrations of cycle-star and melon graphs.
+Observe that any cycle-star (respectively, melon) graph can be fully described by a set of integers specifying the number of vertices on each cycle (respectively, path).
 This representation allows us to relate such graphs directly to integer sets, a connection that underlies both our hardness proofs and our algorithmic techniques.
-These graph families are also simple representatives of well-studied classes: both have bandwidth 2, $k$-cycle graphs form a subclass of cactus graphs, and $k$-path graphs form a subclass of series–parallel graphs.
-Consequently, our hardness results extend beyond $k$-cycles and $k$-paths to broader sparse families, including series–parallel graphs, for which no hardness results were previously known.
-Definition 2.3 (Graph Bandwidth).
-Given a graph $G$, an ordering with bandwidth $k$ of $G$ is a permutation $v_{1},\ldots,v_{n}$ of vertices of $G$ such that for each edge $v_{i}\sim v_{j}$ we have $|i-j|\leq k$.
-The bandwidth of $G$ is the minimum $k$ such that such an ordering exists.
-Certain graphs of bounded bandwidth have also been studied in the context of telephone broadcasting.
-For instance, necklace graphs—collections of cycles where each consecutive pair shares a vertex—have bandwidth $k=2$ and the problem can be optimally solved for them [13, 15].
+These graph families are also simple representatives of well-studied classes: cycle-star graphs form a subclass of cactus graphs, and melon graphs form a subclass of series–parallel graphs.
+Consequently, our hardness results extend beyond cycle-star and melon graphs to broader sparse families, including series–parallel graphs, for which no hardness results were previously known.
+Definition 2.3 (Graph Cutwidth).
+The cutwidth of a graph $G=(V,E)$, denoted $\mathrm{cw}(G)$, is the minimum integer $k$ such that there exists a linear ordering $\pi=(v_{1},\ldots,v_{n})$ of $V$ with the property that, for every $i\in[n]$, at most $k$ edges have one endpoint in $\{v_{1},\ldots,v_{i}\}$ and the other in $\{v_{i+1},\ldots,v_{n}\}$.
+Computing the cutwidth is known to be fixed-parameter tractable.
+In particular, for every fixed constant $k$, there exists a linear-time algorithm that either constructs a cutwidth-$k$ vertex ordering or correctly reports that the cutwidth exceeds $k$ [28].
+This makes cutwidth a particularly well-suited parameter for algorithmic approaches based on linear layouts and dynamic programming, such as those developed in this paper.
+Definition 2.4 (Split graphs).
+A graph $G=(V,E)$ is a split graph if its vertex set can be partitioned into two sets $C$ and $I$ such that the subgraph induced by $C$ is a clique and the subgraph induced by $I$ is an independent set.
 3 Hardness Results
-In this section, we present our hardness results for Telephone Broadcasting in $k$-cycle graphs and $k$-path graphs.
+In this section, we present our hardness results for Telephone Broadcasting in cycle-star graphs and melon graphs.
 These results extend the existing hardness landscape by showing that even some of the simplest graph families exhibit computational intractability.
-Recent hardness results for broadcasting typically rely on reductions from variants of the Numerical Matching with Target Sums (NMTS) problem, which is known to be strongly NP-hard [11] (Problem SP17).
+Recent hardness results for Telephone Broadcasting typically rely on reductions from variants of the Numerical Matching with Target Sums (NMTS) problem, which is known to be strongly NP-hard (Problem SP17 of  [13]).
 An instance of NMTS consists of three multisets $U,V,$ and $W$, each containing $m$ positive integers.
@@
 [8] studied another restricted variant of NMTS, in which all members of $U$ are even while all members of $V\cup W$ are odd.
-They proved that this variant remains strongly NP-hard and provided a reduction to broadcasting in graphs that are distance-1 from a path forest.
+They proved that this variant remains strongly NP-hard and provided a reduction to Telephone Broadcasting in graphs that are distance-1 from a path forest.
 To establish our hardness results, we rely on another restricted version of NMTS in which $U=V=[m]$.
@@
 Definition 3.1 (Restricted Numerical 3-Dimensional Matching (RN3DM) [29]).
-Given a multiset $W=\{w_{1},w_{2},\ldots,w_{m}\}$ of positive integers such that $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$, do there exist two permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)+w_{i}=e\quad\text{for all }1\leq i\leq m\,$?
+Given a multiset $W=\{w_{1},w_{2},\ldots,w_{m}\}$ of positive integers such that $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$, do there exist two permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)+w_{i}=e\quad\text{for all }1\leq i\leq m\,$?
 Example 1 (Yes-instance of RN3DM).
 Consider the instance of RN3DM with $W=\{1,3,4,4\}$, so $m=4$.
-Note that we have $\sum_{i=1}^{m}w_{i}+m(m+1)=32=me$ for $e=8$.
+Note that we have $\sum_{i=1}^{m}w_{i}+m(m+1)=32=m\cdot e$ for $e=8$.
 Let $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$.
@@
 Given that RN3DM is strongly NP-hard [29], we assume that the numbers in $W$ are encoded in unary.
-We present reductions from RN3DM to broadcasting in $k$-cycle and $k$-path graphs.
-For $k$-cycle graphs, our approach introduces an intermediate problem as part of the reduction series, whereas for $k$-path graphs, we directly reduce RN3DM to telephone broadcasting.
-3.1 Hardness of Telephone Broadcasting in $k$-cycle graphs
-We prove that the $k$-cycle broadcasting decision problem is NP-hard through a series of reductions starting from RN3DM.
+We present reductions from RN3DM to Telephone Broadcasting in cycle-star and melon graphs.
+For cycle-star graphs, our approach introduces an intermediate problem as part of the reduction series, whereas for melon graphs, we directly reduce RN3DM to Telephone Broadcasting.
+3.1 Hardness of Telephone Broadcasting in Cycle-star Graphs
+We prove that the cycle-star Telephone Broadcasting is NP-hard through a series of reductions starting from RN3DM.
 As a first step, we define an intermediate problem, called Even-Odd RN3DM, which is a variant of RN3DM.
@@
 Definition 3.2 (Even-Odd Restricted Numerical 3-Dimensional Matching (Even-Odd RN3DM)).
-Given a multiset $C=\{c_{1},c_{2},\ldots,c_{m}\}$ of positive integers such that $\sum_{i=1}^{m}c_{i}=m(2m+1),$ does there exist a permutation $\alpha$ of $2,4,6,\ldots,2m$} and a permutation $\beta$ of $1,3,5,\ldots,2m-1$} such that $\alpha(i)+\beta(i)=c_{i}\quad\text{for all }1\leq i\leq m\,?$
+Given a multiset $C=\{c_{1},c_{2},\ldots,c_{m}\}$ of positive integers such that $\sum_{i=1}^{m}c_{i}=m(2m+1),$ does there exist a permutation $\alpha$ of $\{2,4,6,\ldots,2m\}$ and a permutation $\beta$ of $\{1,3,5,\ldots,2m-1\}$ such that $\alpha(i)+\beta(i)=c_{i}\quad\text{for all }1\leq i\leq m\,?$
 Observe that any instance of Even-Odd RN3DM in which $C$ contains an even element is trivially a No-instance, since each $c_{i}$ must be represented as the sum of one even and one odd number, which always yields an odd value.
@@
 One can check for every $i\in[4]$, $\alpha(i)+\beta(i)=c_{i}$.
-For instance $\alpha(1)+\beta(1)=4+3=7=c_{1}$, and $\alpha(2)+\beta(2)=6+1=7=c_{2}$.
+For instance $\alpha(1)+\beta(1)=4+3=c_{1}$, and $\alpha(2)+\beta(2)=6+1=c_{2}$.
 The same holds for $i=3,4$.
@@
 […]
-Figure 2: An instance $C=\{7,7,9,13\}$ of Even-Odd RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=8$).
-The broadcast tree associated with with solution $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$ is highlighted.
+Figure 2: An instance $C=\{7,7,9,13\}$ of Even-Odd RN3DM reduces to the above instance of Telephone Broadcasting in cycle-star graphs (with $s$ as the originator and $br=8$).
+The broadcast tree associated with solution $\alpha=(4,6,2,8)$ and $\beta=(3,1,7,5)$ is highlighted.
 Lemma 3.3.
@@
 Given an instance $W=\{w_{1},\ldots,w_{m}\}$ of RN3DM with target sum $e$, we construct an instance $C=\{c_{1},\ldots,c_{m}\}$ of Even-Odd RN3DM by setting $c_{i}=2e-2w_{i}-1\quad\text{for each }i\in[m].$ It is straightforward to verify that $W$ is a Yes-instance of RN3DM if and only if $C$ is a Yes-instance of Even-Odd RN3DM.
-The transformation uses unary encodings, so the reduction is strong.
 $\Box$
@@
 Suppose $C$ is a yes-instance of Even-Odd RN3DM.
-Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that for each $i\in[m]$, $\alpha(i)+\beta(i)=c_{i}.$ Define $\lambda(i)=\alpha(i)/2$ and $\mu(i)=(\beta(i)+1)/2$.
+Then there exist permutations $\alpha$ of $\{2,4,\ldots,2m\}$ and $\beta$ of $\{1,3,\ldots,2m-1\}$ such that for each $i\in[m]$, $\alpha(i)+\beta(i)=c_{i}.$ Define $\lambda(i)=\alpha(i)/2$ and $\mu(i)=(\beta(i)+1)/2$.
 Both $\lambda$ and $\mu$ are permutations of $[m]$.
@@
 $\lambda(i)+\mu(i)+w_{i}=\frac{\alpha(i)}{2}+\frac{\beta(i)+1}{2}+w_{i}=\frac{c_{i}+1}{2}+w_{i}=\frac{(2e-2w_{i}-1)+1}{2}+w_{i}=e.$
-Thus $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
+Thus, $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
 Since the reduction preserves yes-instances in both directions and can be carried out in polynomial time, Even-Odd RN3DM is strongly NP-hard. $\Box$
+We are ready to prove the main result of this section, namely, the hardness of Telephone Broadcasting in cycle-star.
 Theorem 3.4.
-Broadcasting in $k$-cycle graphs is NP-complete.
+Telephone Broadcasting in cycle-star graphs is NP-Complete.
 Proof.
-Telephone Broadcasting is in NP in general graphs (and hence also for $k$-cycle graphs).
+Telephone Broadcasting is in NP in general graphs (and hence also for cycle-star graphs).
 We prove NP-hardness by a reduction from the Even-Odd RN3DM problem, which is strongly NP-hard by Lemma 3.3.
 Let $C=\{c_{1},c_{2},\ldots,c_{m}\}$ be an instance of Even-Odd RN3DM.
-We construct an instance $(G,s,t)$ of Telephone Broadcasting in $k$-cycle graphs as follows.
+We construct an instance $(G,s,t)$ of Telephone Broadcasting in cycle-star graphs as follows.
 Graph $G$ consists of $m$ cycles, containing $c_{1},c_{2},\ldots,c_{m}$, extra vertices in addition to a single vertex $s$ that is shared between them.
@@
 Suppose the answer to instance $C$ of Even-Odd RN3DM is yes.
-Then there exist permutations $\alpha$ of $(2,4,\ldots,2m)$ and $\beta$ of $(1,3,\ldots,2m-1)$ such that $\alpha(i)+\beta(i)=c_{i}$ for all $i\in[m]$.
+Then there exist permutations $\alpha$ of $\{2,4,\ldots,2m\}$ and $\beta$ of $\{1,3,\ldots,2m-1\}$ such that $\alpha(i)+\beta(i)=c_{i}$ for all $i\in[m]$.
 Construct a broadcast scheme as follows: the source $s$ informs its two neighbors in each cycle $c_{i}$ at times $2m-\alpha(i)+1$ and $2m-\beta(i)+1$, respectively.
@@
 Each newly informed vertex continues by informing its uninformed neighbor in the subsequent round.
-Hence, by round $br=2m$, exactly $\alpha(i)+\beta(i)=c_{i}$ vertices of the cycle (excluding $s$) are informed, and therefore the entire graph is informed.
-Thus, $(G,s,br)$ is a yes-instance of broadcasting.
+Hence, by round $br=2m$, exactly $\alpha(i)+\beta(i)=c_{i}$ vertices of the cycle (excluding $s$) are informed, and, therefore, the entire graph is informed.
+Thus, $(G,s,br)$ is a yes-instance of Telephone Broadcasting.
 (Reverse direction).
-Now suppose $(G,s,br)$ is a yes-instance of broadcasting with $br=2m$.
-So, there is a broadcast scheme that completes within $22m$ rounds.
+Now suppose $(G,s,br)$ is a yes-instance of Telephone Broadcasting with $br=2m$.
+Therefore, there is a broadcast scheme that completes within $2m$ rounds.
 Without loss of generality, assume the scheme is non-lazy, i.e., every informed vertex calls an uninformed neighbor whenever possible.
@@
 $\Box$
-3.2 Hardness of Telephone Broadcasting in $k$-Path Graphs
-Next, we use a reduction from RN3DM to establish NP-completeness of broadcasting in $k$-path graphs.
+3.2 Hardness of Telephone Broadcasting in Melon Graphs
+Next, we use a reduction from RN3DM to establish NP-Completeness of Telephone Broadcasting in melon graphs.
 Theorem 3.5.
-Broadcasting in $k$-path graphs is NP-complete.
+Telephone Broadcasting in melon graphs is NP-Complete.
+Proof sketch: We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-hard [29].
+Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM.
+By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$.
+We construct an instance $(G,s,br)$ of Telephone Broadcasting in melon graphs as follows.
+The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$.
+Furthermore, we add an edge connecting $s$ and $t$.
+The vertex $s$ serves as the source of the broadcast.
+The total number of vertices in $G$ is $n=2+\sum_{i=1}^{m}(e-w_{i})=2+me-\sum_{i=1}^{m}w_{i}=2+m(m+1).$ Both $s$ and $t$ have degree $2m+1$.
+Since RN3DM is strongly NP-hard, we may assume that the values $w_{i}$ are encoded in unary, ensuring that the reduction is polynomial time.
+The decision problem asks whether broadcasting completes within $br=m+1$ rounds (see Figure 3).
+The answer to this decision problem is yes if and only if the $W$ is a yes-instance of RN3DM.
+$\Box$
 Proof.
-Telephone Broadcasting belongs to NP for general graphs and hence also for $k$-path graphs.
+Telephone Broadcasting belongs to NP for general graphs and hence also for melon graphs.
 We prove NP-hardness by a reduction from the Restricted Numerical 3-Dimensional Matching (RN3DM) problem (Definition 3.1), which is strongly NP-hard [29].
 Let $W=\{w_{1},w_{2},\ldots,w_{m}\}$ be an instance of RN3DM.
-By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$.
-We construct an instance $(G,s,br)$ of Telephone Broadcasting in $k$-path graphs as follows.
+By definition, $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$.
+We construct an instance $(G,s,br)$ of Telephone Broadcasting in melon graphs as follows.
 The graph $G$ consists of $m$ internally disjoint paths with shared endpoints $s$ and $t$, where the $i$-th path contains $\ell_{i}=e-w_{i}$ internal vertices in addition to $s$ and $t$.
@@
 […]
-Figure 3: An instance $W=\{1,3,4,4\}$ of RN3DM reduces to the above instance of telephone broadcasting in $k$-cylce graphs (with $s$ as the originator and $br=5$).
+Figure 3: An instance $W=\{1,3,4,4\}$ of RN3DM reduces to the above instance of Telephone Broadcasting in cycle-star graphs (with $s$ as the originator and $br=5$).
 The broadcast tree associated with with solution $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$ is highlighted.
-4 Polynomial Time Approximation Schemes
-4.1 PTAS for $k$-Cycle Graphs
-In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-cycle graphs.
-Recall that a $k$-cycle graph is obtained by joining $k$ cycles at a common central vertex $r$.
-We assume that the number of cycles kk—and hence the optimal broadcast time—is asymptotically large; for constant $k$, the problem can be solved by a straightforward exhaustive search (see, e.g., [4]).
-In what follows, we design a PTAS for the case where the broadcast originates at $s=r$.
+4 Efficient Polynomial Time Approximation Schemes
+4.1 EPTAS for Cycle-star Graphs
+In this section, we present an efficient polynomial-time approximation scheme (EPTAS) for Telephone Broadcasting in cycle-star graphs.
+Recall that a cycle-star graph is obtained by joining $k$ cycles at a common central vertex $r$.
+We assume that the number of cycles kk—and hence the optimal broadcast time—is asymptotically large; for constant $k$, the problem can be solved optimally by a straightforward exhaustive search (see, e.g., [5]).
+In what follows, we design an EPTAS for the case where the broadcast originates at $s=r$.
 This restriction is without loss of generality: any $c$-approximation for the case that the center is the originator (when $s=r$) can be extended to a $c$-approximation in the general case ($s\neq r$).
-Suppose the originator $s$ is a non-center vertex on some cycle $C_{m}$, at distance $d\geq 1$ from $r$.
+Suppose the originator $s$ is a non-center vertex on some cycle $Y$, at distance $d\geq 1$ from $r$.
 The message can first be forwarded along the shortest path from $s$ to $r$.
-Once $r$ is informed, it first sends the message to its other neighbor on $C_{m}$ and then proceeds with the $c$-approximation algorithm on the rest of the graph.
-If the overall broadcast time is determined by a vertex in $C_{m}$, then the scheme is optimal.
-Otherwise, the process completes within $d+1+c\cdot\boldsymbol{br}^{-}$ rounds, where $\boldsymbol{br}^{-}$ is the optimal broadcast time excluding $C_{m}$.
+Once $r$ is informed, it first sends the message to its other neighbor on $Y$ and then proceeds with the $c$-approximation algorithm on the rest of the graph.
+If the overall broadcast time is determined by a vertex in $Y$, then the scheme is optimal.
+Otherwise, the process completes within $d+1+c\cdot\boldsymbol{br}^{-}$ rounds, where $\boldsymbol{br}^{-}$ is the optimal broadcast time excluding $Y$.
 Since ${\boldsymbol{br}^{*}}\geq d+\boldsymbol{br}^{-}$, this yields the same approximation factor $c$.
-We introduce the following problem, which is the key part of our PTAS algorithms.
+From Telephone Broadcasting in Cycle-star Graphs to Prefix Covering
+We introduce the following problem, which is the key part of our EPTAS algorithms.
 Definition 4.1.
 The input to the Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers.
-The objective is to determine the minimum integer $m$ such that the set $C=[m]$ can be partitioned into $n$ subsets $C_{0},C_{1},\ldots,C_{n}$, such that for each $i\in[n]$, we have $|C_{i}|\leq 2$ and $\sum_{x\in C_{i}}x\,\geq\,\ell_{i}$.
+The objective is to determine the minimum integer $m$ such that the set $C=[m]$ can be partitioned into $n+1$ subsets $C_{0},C_{1},\ldots,C_{n}$, such that for each $i\in[n]$, we have $|C_{i}|\leq 2$ and $\sum_{x\in C_{i}}x\,\geq\,\ell_{i}$.
 In this context, we say that $\ell_{i}$ is covered by the elements of $C_{i}$, and we refer to $C$ as the covering set of $S$.
@@
 When $S=\{13,8,7,6\}$, the optimal solution is given by $C=[8]$, and we have $C_{1}=\{8,5\},C_{2}=\{7,2\},C_{3}=\{4,3\},C_{4}=\{6\}$ (and $C_{0}=\{1\}$).
-We begin by establishing the equivalence between Telephone Broadcasting in $k$-cycle graphs and the Prefix Covering problem.
+We begin by establishing the equivalence between Telephone Broadcasting in cycle-star graphs and the Prefix Covering problem.
 Lemma 4.2.
-Let $G$ be a $k$-cycle graph with central vertex $s$, where the $i$-th cycle contains $c_{i}$ additional vertices besides $s$, and let $S=\{c_{1},\ldots,c_{k}\}$.
+Let $G$ be a cycle-star graph with central vertex $s$, where the $i$-th cycle contains $c_{i}$ additional vertices besides $s$, and let $S=\{c_{1},\ldots,c_{k}\}$.
 Then Telephone Broadcasting for instance $(G,s)$ becomes equivalent to Prefix Covering for $S$.
@@
 Thus, every covering for the Prefix Covering instance corresponds to a valid broadcast scheme with $\boldsymbol{br}$ rounds.
-Therefore, Telephone Broadcasting in $k$-cycle graphs is equivalent to the Prefix Covering problem. $\Box$
-4.1.1 PTAS for Prefix Covering
-In this section, we present a PTAS for the Prefix Covering problem, that is, an algorithm with approximation factor $(1+\epsilon)$ for a given $\epsilon>1$.
-By Lemma 4.2, this would give a PTAS for Telephone Broadcasting in $k$-cycle graphs.
+Therefore, Telephone Broadcasting in cycle-star graphs is equivalent to the Prefix Covering problem. $\Box$
+EPTAS for Prefix Covering
+In this section, we present an EPTAS for the Prefix Covering problem, that is, an algorithm with approximation factor $(1+\epsilon)$ for a given $\epsilon>0$.
+By Lemma 4.2, this would give an EPTAS for Telephone Broadcasting in cycle-star graphs.
 As a high-level view of the algorithm, it addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity.
-The main idea is to transform the original ground multiset and the covering set into a simplified version with constant support (distinct elements), making it feasible to apply an exhaustive search which maintains provable approximation guarantees.
+The main idea is to transform the original ground multiset and the covering set into a simplified version with constant support (i.e., distinct elements), enabling formulation as a feasibility ILP with a constant number of variables.
+Solving this ILP yields to the desired EPTAS.
 The algorithm operates as follows.
@@
 Hence, verifying whether $\mathfrak{R}_{p}(C)$ successfully covers $\mathfrak{R}_{p}(S)$ is sufficient for approximating the original instance.
-This verification can be performed by exhaustive search over $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$.
+This verification can be performed by solving an ILP over $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$.
 Since both sets have constant support (bounded by $p$), this step can be carried out in time independent of $n$ (Lemma 4.5).
@@
 Lemma 4.5.
-Given two sets $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering set $\mathfrak{R}_{p}(C)$ (if one exists) or correctly reports that no such covering is possible.
+Given two sets $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, there exists an algorithm that, in time $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ (if one exists) or correctly reports that no such covering is possible.
+Proof sketch: We model the covering problem as a feasibility instance of integer linear programming with one variable for each possible way of covering a distinct element in $\mathfrak{R}_{p}(S)$ using one or two elements in $\mathfrak{R}_{p}(C)$.
+Given that both the ground multiset $\mathfrak{R}_{p}(S)$ and the covering multiset $\mathfrak{R}_{p}(C)$ have constant support $p$, there are only $\mathcal{O}(p^{3})$ such variables.
+The constraints enforce that (i) every ground element is covered exactly once, (ii) no covering type is used more than its multiplicity, and (iii) only feasible type combinations are allowed.
+By Lenstra’s algorithm, the resulting integer linear feasibility can be solved in time $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$.
+Thus, we can either find a feasible covering or correctly conclude that none exists.
+$\Box$
 Proof.
-Both the ground set $\mathfrak{R}_{p}(S)$ and the covering set $\mathfrak{R}_{p}(C)$ have constant support $p$.
-Fix a distinct value $v\in\mathfrak{R}_{p}(S)$.
-For each pair $(a,b)$ of elements from $\mathfrak{R}_{p}(C)$, we can encode how many times $(a,b)$ is used to cover elements of $\mathfrak{R}_{p}(S)$ equal to $v$.
-Since there are at most $|S|$ such elements, this count can be represented in $\log|S|$ bits.
-As there are at most $p^{2}$ pairs from $\mathfrak{R}_{p}(C)$ and $p$ distinct values in $\mathfrak{R}_{p}(S)$, any candidate covering can be encoded using at most $p^{3}\log|S|$ bits.
-Hence, the number of candidate solutions is bounded by $\mathcal{O}(2^{p^{3}}|S|)$.
-Checking whether a candidate encoding yields a valid covering takes $\mathcal{O}(|S|+|C|)$ time.
-Therefore, the total running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+|C|))$.
-Moreover, whenever a valid covering exists, the algorithm outputs it by interpreting the corresponding encoding; otherwise, it correctly reports that no covering is possible.
+Let $V$ and $T$ be the set of distinct elements in $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, respectively.
+Then $|V|\leq p$ and $|T|\leq p$.
+For each pair $\{i,j\}$ of items in $T$ and each ground type $v\in V$, introduce an integer variable $x_{i,j,v}\in\mathbb{Z}_{\geq 0}$, intended to denote how many times the pair $(i,j)$ is used to cover elements of type $v$.
+Similarly, for each element $i\in T$ and ground type $v\in V$, we let $x_{i,v}\in\mathbb{Z}_{\geq 0}$ to denote the number of times $i$ is used to cover elements of type $v$.
+There are $\big({{|T|}\choose{2}}+|T|\big)|V|\leq p^{3}$ variables.
+We write a feasibility ILP whose constraints enforce that (i) every ground element is covered exactly once, and (ii) cover-items are not used more than their multiplicities.
+Cover all ground elements.
+For each $v\in V$, we let $\mu_{S}(v)$ be the multiplicity of $v$, that is, $\mu_{S}(v):=|\{s\in\mathfrak{R}_{p}(S):s=v\}|$.
+To ensure $v$ is covered exactly once, we add a constraint:
+$\sum_{i,j\in T}x_{i,j,v}+\sum_{i\in T}x_{i,v}\;=\;\mu_{S}(v).$ (1)
+Do not overuse cover-items.
+For each $t\in T$, let $\mu_{C}(t)$ be the multiplicity of type $t$ in $\mathfrak{R}_{p}(C)$, that is, $\mu_{C}(t):=|\{s\in\mathfrak{R}_{p}(C):s=t\}|$.
+Each time we use a pair $(i,j)$ we consume one item of type $i$ and one item of type $j$ (with the understanding that if $i=j$ then we consume two copies).
+Thus, for each $t\in T$ we impose
+$\sum_{v\in V}\ \big(\sum_{i,j\in T}\bigl(\mathbf{1}[i=t]+\mathbf{1}[j=t]\bigr)\,x_{i,j,v}+\sum_{i\in T}\bigl(\mathbf{1}[i=t]\big)\,x_{i,v}\bigr)\;\leq\;\mu_{C}(t).$ (2)
+Here, $\mathbf{1}[j=t]$ denotes the indicator function, which is equal to $1$ if $j=t$ and 0 otherwise.
+Type-feasibility.
+Finally, we forbid pairs that cannot cover type $v$ by setting $x_{i,j,v}=0$ whenever $(i,j)$ is not a valid set for covering $v$, that is, $i+j<v$.
+Similarly, we force $x_{i,v}=0$ when $i<v$.
+This ILP has at most $p^{3}$ integer variables, and the number of constraints is $\mathcal{O}(p^{3})$.
+By Lenstra’s theorem [22] (and subsequent improvements by Kannan [21]), integer linear feasibility with a fixed number $d$ of variables can be solved in time $d^{\mathcal{O}(d)}\cdot\operatorname{poly}(L)$, where $L$ is the encoding length of the ILP.
+Here $d\leq p^{3}$ and $L=\operatorname{poly}(|S|+|C|+p)$, hence the feasibility test runs in $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$ time.
+If feasible, the resulting integer solution specifies a valid covering; otherwise, no such solution exists.
 $\Box$
-To summarize, given a covering guess size $m$ and ground set $\mathfrak{R}_{p}(S)$, we form $\mathfrak{R}_{p}(C)$, and check whether we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ with an exhaustive approach (Lemma 4.5).
+To summarize, given a covering guess size $m$ and ground set $\mathfrak{R}_{p}(S)$, we form $\mathfrak{R}_{p}(C)$, and check whether we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ with a feasibility ILP (Lemma 4.5).
 If we cannot find a covering, we return bad-guess; we cannot cover $\mathfrak{R}_{p}(S)$ using $[m]$.
@@
 Applying the Proposition 4.4, we can get an approximating factor of $(1+2/p)(1+1/p)$, which can be as close as we want to $1$ by setting $p$ large enough.
-Using binary search adds a multiplicative factor of $\log\boldsymbol{OPT}$ to the runtime we used for the exhaustive search (see Lemma 4.5).
-Hence, the overall running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$.
+Using binary search adds a multiplicative factor of $\log\boldsymbol{OPT}$ to the runtime we used for ILP (see Lemma 4.5).
+Hence, the overall running time is $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$.
 We can conclude the following:
 Theorem 4.6.
-There is a PTAS for the Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
+There is a $(1+\epsilon)$-approximation for the Prefix Covering problem that terminates in $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
 Proof.
@@
   2. (ii).
-    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
+    Use the feasibility ILP (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
 Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$.
@@
 Using Proposition 4.4, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)(1+2/p)\boldsymbol{OPT}(S)$.
-The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$.
+The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$ time, giving total complexity $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$.
 $\Box$
-Lemma 4.2 and Theorem 4.6 imply the following Corollary.
+EPTAS for Telephone Broadcasting in Cycle-Star graphs
+Given the equivalence between Telephone Broadcasting in cycle-star graphs and the Prefix Covering problem, as established in Lemma 4.2, along with the EPTAS presented in Theorem 4.6 for the Prefix Covering problem, we can conclude the following.
 Corollary 4.7.
-There exists a PTAS for Telephone Broadcasting in $k$-cycle graphs.
-In particular, for a $k$-cycle graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
-4.2 PTAS for $k$-Path Graphs
-In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-path graphs. Consider a $k$-path graph $G$ with endpoints $s$ and $t$.
+There exists an EPTAS for Telephone Broadcasting in cycle-star graphs.
+In particular, for a cycle-star graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
+Extension to multi-source broadcasting
+The EPTAS of Corollary 4.7 extends naturally to the multi-source broadcasting setting. First, we provide a sketch of this extension.
+The main idea is to guess the optimal broadcast time $\boldsymbol{br}$ and, for each guess, attempt to construct a broadcast schedule.
+If $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, the procedure returns a schedule of length at most $(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1$.
+Trying all relevant guesses for $\boldsymbol{br}$ and selecting the shortest resulting schedule yields an EPTAS.
+For each fixed guess $\boldsymbol{br}$, we reduce the multi-source instance to a single-source instance as follows.
+We first perform a preprocessing step that removes a window of at most $\boldsymbol{br}$ vertices around each non-center source.
+These vertices can always be informed within $\boldsymbol{br}+1$ rounds directly from their respective sources.
+If this preprocessing disconnects the graph or leaves a vertex at a distance larger than $\boldsymbol{br}$ from every source, we reject the guess value $\boldsymbol{br}$.
+Otherwise, the remaining graph $G^{\prime}$ is a graph of distance 1 to a path forest (See Figure 1).
+This graph has the center as the unique bottleneck through which all remaining vertices must be informed.
+We therefore reduce the problem to a single-source instance $(G^{\prime},s)$, where $s$ is the center.
+This instance is transformed into a modified version of the Prefix Covering problem in which cycle components allow one or two-element covers as before, while path components (created by preprocessing) correspond to ground elements that must be covered by a single item.
+The rounding and approximation steps from the single-source EPTAS remain unchanged, and the resulting rounded instance of Prefix Covering has constant support.
+It can therefore be solved exactly using a fixed-dimension ILP as before.
+Now, we present details on the extension of the EPTAS of Corollary 4.7 to the multi-source broadcasting setting.
+Suppose we are given an instance $(G,S)$, where $S$ is the set of all sources, and a fixed parameter $\epsilon>0$.
+Let $\boldsymbol{br}$ be a broadcast-time guess value.
+We describe a procedure that, when $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, outputs a valid broadcasting scheme that completes within $(1+\epsilon)\boldsymbol{br}+1$.
+Otherwise, it returns either Bad-Guess or a scheme with no guarantee on its broadcast time.
+As we do not know the optimal broadcast time, we iterate over all $\boldsymbol{br}\in[\lceil\log n\rceil,n]$ and take the scheme with the best broadcasting time among all returned schemes.
+This way, we can achieve broadcast time $(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1$.
+Preprocessing.
+For each non-center source $s_{i}$, we remove all vertices that lie within distance $\boldsymbol{br}$ of $s_{i}$ on either of the two paths connecting $s_{i}$ to the center, excluding the center itself.
+(The vertex $s_{i}$ is removed as well.) If the center is not initially a source, let $s_{m}$ be the source closest to the center, at distance $d_{\min}$, and remove up to $\boldsymbol{br}-d_{\min}$ additional vertices along the remaining path between the center and $s_{m}$.
+Let $G^{\prime}$ denote the remaining graph.
+If $G^{\prime}$ is disconnected, or if $d_{\min}>\boldsymbol{br}$, then there exists a vertex whose distance from every source exceeds $\boldsymbol{br}$, and we correctly return Bad-Guess.
+Otherwise, $G^{\prime}$ will be a graph of distance $1$ to path forests (see Figure 1).
+The preprocessing removes a window of at most $\boldsymbol{br}$ vertices around each non-center source.
+Note that all removed vertices can be informed within $\boldsymbol{br}+1$ rounds by broadcasting from their respective sources.
+Reduction to covering.
+Consider an instance $(G^{\prime},s)$ of single-source Telephone Broadcasting with the center being the source.
+We associate $(G^{\prime},s)$ with an instance $\mathcal{S}$ of the Prefix Covering problem, defined analogously to Lemma 4.2 for cycle-star graphs.
+The only difference is that some elements of the ground multiset $\mathcal{S}$, which we refer to as path elements, correspond to paths in $G^{\prime}$ rather than cycles.
+Each path element must be covered by a subset of size exactly $1$ (it cannot be covered with two elements from the ground set).
+This modified covering instance can be solved in the same way as discussed in Section 4.1.
+In particular, the rounding steps remain unchanged.
+The only additional constraint arises in the final step (Lemma 4.5), where we ensure that each path element in the rounded ground set is covered by a single-item subset.
+This produces a broadcast schedule for $(G^{\prime},s)$ that completes within $(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)$ rounds.
+Combining the solutions. Overall, broadcasting in the original multi-source instance $G$ completes within
+$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}.$
+The term $\boldsymbol{br}+1$ accounts for informing all vertices removed during the preprocessing step.
+The second term captures the time required to inform the center and its neighbor along the path toward $s_{m}$ (namely, $d_{\min}+1$ rounds), followed by the execution of the single-source broadcast within $G^{\prime}$.
+Suppose $br={\boldsymbol{br}^{*}}(G,S)$.
+Now, if $G^{\prime}$ has no vertices other than the center, then we have found a broadcasting schedule with $\boldsymbol{br}+1={\boldsymbol{br}^{*}}(G,S)+1$ rounds.
+Otherwise, in the optimal scheme, vertices in $G^{\prime}$ must be informed through the center.
+This is because the vertices in $G^{\prime}$ are the ones that were not removed when forming $G^{\prime}$ and thus are at a distance further than $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$ from the sources in their respective cycles.
+Therefore, as the center cannot be informed earlier than round $d_{\min}$, we can write ${\boldsymbol{br}^{*}}(G,S)\geq d_{\min}+{\boldsymbol{br}^{*}}(G^{\prime},s)$.
+Hence, the broadcast time of our scheme is at most
+$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}\leq(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1.$
+4.2 EPTAS for Melon Graphs
+In this section, we present an efficient polynomial-time approximation scheme (EPTAS) for Telephone Broadcasting in melon graphs.
+Consider a melon graph $G$ with endpoints $s$ and $t$.
 In what follows, we assume that $s$ is the designated source, holding the message at time 0, and that $t$ acts as a late source, receiving the message at time $\alpha$, with $\alpha\leq d(s,t)$, where $d(s,t)$ denotes the distance between $s$ and $t$ in $G$.
 We show that any instance of the (single-source) problem with an arbitrary originator can be reduced to this double-source setting (with $s$ as the source and $t$ as the late source) without changing the approximation factor (Lemma 4.8).
-In Lemma 4.10, we establish an equivalence between double-source broadcasting and the integer covering problem that we call Double Prefix Covering (Definition 4.9).
-Finally, by applying a rounding-and-covering approach similar to the one used for Prefix Covering in $k$-cycle graphs, we obtain a PTAS for Double Prefix Covering, and consequently for Telephone Broadcasting in $k$-path graphs (Section 4.2.1).
+In Lemma 4.10, we establish an equivalence between double-source Telephone Broadcasting and the integer covering problem that we call Double Prefix Covering (Definition 4.9).
+Finally, by applying a rounding-and-covering approach similar to the one used for Prefix Covering in cycle-star graphs, we obtain an EPTAS for Double Prefix Covering, and consequently for Telephone Broadcasting in melon graphs (Section 4.2).
 In what follows, we describe these steps in detail.
-First, we prove the following lemma, which implies we can restrict our attention to the double-source broadcasting.
+First, we prove the following lemma, which implies we can restrict our attention to the double-source Telephone Broadcasting.
 Lemma 4.8.
-Any $c$-approximation algorithm for Telephone Broadcasting in $k$-path graphs in the double-source setting—where $s$ is the primary source and $t$ is a late source receiving the message at time $\alpha$—can be extended to a $c$-approximation algorithm for the (single-source) general setting, where the originator $v_{o}$ may be any vertex of the graph.
+Any $c$-approximation algorithm for Telephone Broadcasting in melon graphs in the double-source setting—where $s$ is the primary source and $t$ is a late source receiving the message at time $\alpha$—can be extended to a $c$-approximation algorithm for the (single-source) general setting, where the originator $v_{o}$ may be any vertex of the graph.
 Proof sketch: Without loss of generality, assume $v_{o}$ is closer to $s$ than to $t$.
@@
 Without loss of generality, assume $d(v_{o},s)\leq d(v_{o},t)$.
-Let $d(v_{o},s)$.
 The message is first forwarded along the shortest path from $v_{o}$ to $s$, which takes $d(v_{o},s)$ rounds.
@@
 $\Box$
-We introduce the following problem, which is the key part of our PTAS algorithms.
+From Telephone Broadcasting in Melon Graphs to Double Prefix Covering
+We introduce the following problem, which is the key part of our EPTAS.
 Definition 4.9.
@@
 The objective is to determine the minimum integer $m$ such that there exist functions $c:S\to\{0\}\cup[m]$ and $d:S\to\{0\}\cup[m-\beta]$ satisfying $c(i)+d(i)\geq\ell_{i}$ for all $i\in[n]$.
-Moreover, injectivity is required in the following sense: for any $i\neq j$, if both $c(i)>0$ (respectively $d(i)>0$), then $c(i)\neq c(j)$ (respectively $d(i)\neq d(j)$).
+Moreover, injectivity is required in the following sense: for any $i\neq j$, if both $c(i)>0$ (respectively, $d(i)>0$), then $c(i)\neq c(j)$ (respectively $d(i)\neq d(j)$).
 In this setting, we say that each $\ell_{i}$ is covered by the pair $\{c(i),d(i)\}$, and we refer to $C=[m]$ and $D=[m-\beta]$ as the covering sets of $S$.
 Example 8.
-When $S=\{8,4,4,2\},$ the optimal solution is given by $C=[5]$ and $\beta=2$, and we have $c(1)=5,d(1)=3$; $c(2)=4,d(2)=0$, $c(3)=2,d(3)=2$, $c(4)=1,d(4)=1$.
-In what follows, we establish the equivalence between double-source broadcasting in $k$-path graphs and Double Prefix Covering (Definition 4.9).
+When $S=\{8,4,4,2\}$ and $\beta=2$, the optimal solution is given by $C=[5]$, and we have $c(1)=5,d(1)=3$; $c(2)=4,d(2)=0$, $c(3)=2,d(3)=2$, $c(4)=1,d(4)=1$.
+In what follows, we establish the equivalence between double-source Telephone Broadcasting in melon graphs and Double Prefix Covering (Definition 4.9).
 Lemma 4.10.
-Let $G$ be a $k$-path graph formed by $k$ paths with common endpoints $s$ and $t$, where the $i$th path contains $\ell_{i}\geq 0$ internal vertices (in addition to $s$ and $t$).
-Consider a double-source broadcasting instance on $G$, where $s$ receives the message at time 0 and $t$ at time $\alpha$.
+Let $G$ be a melon graph formed by $k$ paths with common endpoints $s$ and $t$, where the $i$th path contains $\ell_{i}\geq 0$ internal vertices (in addition to $s$ and $t$).
+Consider a double-source Telephone Broadcasting instance on $G$, where $s$ receives the message at time 0 and $t$ at time $\alpha$.
 Now, define an instance $I$ of Double Prefix Covering as $I=(S=\{\ell_{1},\ldots,\ell_{k}\},\;\beta=\alpha-1)$.
@@
 $\Box$
-4.2.1 PTAS for Double Prefix Covering
-In the light of Lemmas 4.8 and 4.10, presenting a PTAS for broadcasting in $k$-path graphs reduces to presenting a PTAS for the Double Prefix Covering problem.
+EPTAS for Double Prefix Covering
+In the light of Lemmas 4.8 and 4.10, presenting an EPTAS for Telephone Broadcasting in melon graphs reduces to presenting an EPTAS for the Double Prefix Covering problem.
 This is the focus of the present section.
 Specifically, we prove Theorem 4.13, which establishes the existence of a $(1+\epsilon)$-approximation algorithm for Double Prefix Covering for any $\epsilon>0$.
-Our approach closely parallels the rounding and covering techniques used in Section 4.1.1 for $k$-cycle graphs.
+Our approach closely parallels the rounding and covering techniques used in Section 4.1 for cycle-star graphs.
 In particular, it also addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity.
-The main idea is to transform the original ground multiset and the two covering sets into simplified versions with constant support (i.e., bounded numbers of distinct elements), so that exhaustive search becomes computationally feasible while still preserving provable approximation guarantees.
-Our PTAS for Double Prefix Covering proceeds by reducing the original instance to one with bounded support.
+The main idea is to transform the original ground multiset and the two covering sets into simplified versions with constant support (i.e., bounded numbers of distinct elements), so that an ILP formulation becomes computationally feasible while still preserving provable approximation guarantees.
+Our EPTAS for Double Prefix Covering proceeds by reducing the original instance to one with bounded support.
 Given an input multiset $S$, we first form its rounded version $\mathfrak{R}_{p}(S)$, where each element is rounded up to a larger element of $S$ so that the number of distinct values becomes $p$.
@@
 It therefore suffices to test whether $\mathfrak{R}_{p}(S)$ is coverable using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$.
-Since all three sets have constant support (bounded by $p$), this check can be carried out by exhaustive search in constant time (Lemma 4.12).
+Since all three sets have constant support (bounded by $p$), this check can be carried out by an ILP formulation in polynomial time (Lemma 4.12).
 The remainder of this section provides details of the algorithm.
@@
 This allows us to restrict attention to $\mathfrak{R}_{p}(S)$, which has constant support $p$.
-To cover $\mathfrak{R}_{p}(S)$, we apply a binary strategy.
+To cover $\mathfrak{R}_{p}(S)$, we apply a binary search strategy.
 For a candidate value $m$, we design a procedure that either produces a covering using $m+\lceil m/p\rceil]$, or certifies that no covering exists with $[m]$.
-By iteratively refining $m$, we identify the smallest value for which $\mathfrak{R}_{p}(S)$ can be covered using the rounded covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, where $C=[m]$ and $D=[m-\beta]$.
+This way, we identify the smallest value of $m$ for which $\mathfrak{R}_{p}(S)$ can be covered using the rounded covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, where $C=[m]$ and $D=[m-\beta]$.
 Claim 4.2.
@@
 Lemma 4.12.
-Given three sets $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|+|D|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ or correctly reports that no such covering is possible.
+Given three sets $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, there exists an algorithm that, in time $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|+|D|)$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ or correctly reports that no such covering is possible.
 Proof.
-The ground set $\mathfrak{R}_{p}(S)$ and the covering sets have constant support $p$.
-For any fixed distinct member of $\mathfrak{R}_{p}(S)$ like $v$ and each combination pair of members of $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, we can encode the number of times the pair is used to cover the members of $\mathfrak{R}_{p}(S)$ with value $v$ in $\log|S|$ bits.
-This is because we have at most $|S|$ members with value equal to $v$.
-Given that there are $p^{2}$ distinct pairs from the covering set and $p$ types of value in $\mathfrak{R}_{p}(S)$, any potential solution for the problem can be encoded in $p^{2}.p.\log|S|=p^{3}\log|S|$ bits; we can exhaustively try all and answer whether $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ by trying all potential solutions.
-Thus, the algorithm tries $\mathcal{O}(2^{p^{3}}|S|)$ solutions.
-And, each solution takes linear time to check.
-Therefore, overall it takes $(2^{p^{3}}|S|(|S|+|C|+|D|))$ to decide whether it is possible to cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$.
+Let $V$, $T_{C}$, and $T_{D}$ be the set of distinct elements in $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, respectively.
+Then $|V|\leq p$ and $|T_{C}|,|T_{D}|\leq p$.
+For each tuple $(i,j)\in(T_{C}\cup\{\emptyset\})\times(T_{D}\cup\{\emptyset\})$ (except $(\emptyset,\emptyset)$) and each ground type $v\in V$, introduce an integer variable $x_{i,j,v}\in\mathbb{Z}_{\geq 0}$, intended to denote how many times the pair $(i,j)$ is used to cover elements of type $v$.
+There are $((|T_{C}|+1)(|T_{D}|+1)-1)|V|\leq p^{3}$ variables.
+We write a feasibility ILP whose constraints enforce that (i) every ground element is covered exactly once, and (ii) cover-items are not used more than their multiplicities, (iii) every ground element is covered with the right set of values.
+Cover all ground elements.
+For each $v\in V$, we let $\mu_{S}(v)$ be the multiplicity of $v$, that is, $\mu_{S}(v):=|\{s\in\mathfrak{R}_{p}(S):s=v\}|$.
+To ensure $v$ is covered exactly once, we add a constraint:
+$\sum_{(i,j)\in(T_{C}\cup\{\emptyset\})\times(T_{D}\cup\{\emptyset\})}x_{i,j,v}\;=\;\mu_{S}(v).$ (3)
+Do not overuse cover-items.
+For each $t\in T_{C}$, let $\mu_{C}(t)$ be the multiplicity of type $t$ in $\mathfrak{R}_{p}(C)$, that is, $\mu_{C}(t):=|\{s\in\mathfrak{R}_{p}(C):s=t\}|$.
+Similarly, define $\mu_{D}(t)$ for each $t\in T_{D}$.
+Each time we use a pair $(i,j)$ we consume one item of type $i$ and one item of type $j$ (with the understanding that if $i=j$ then we consume two copies).
+Thus, for each $t\in T_{C}$ we impose
+$\sum_{v\in V}\ \sum_{j\in T_{D}\cup\{\emptyset\}}\,x_{t,j,v}\;\leq\;\mu_{C}(t).$ (4)
+Similarly, we impose the same constraint for $\mu_{D}(t)$ for each $t\in T_{D}$:
+$\sum_{v\in V}\ \sum_{i\in T_{C}\cup\{\emptyset\}}\,x_{i,t,v}\;\leq\;\mu_{D}(t).$ (5)
+Type-feasibility.
+Finally, we forbid pairs that cannot cover type $v$ by setting $x_{i,j,v}=0$ whenever $(i,j)$ is not a valid set for covering $v$, that is, when the sum of $i$ and $j$ (if it exists) is less than $v$.
+This ILP has at most $p^{3}$ integer variables, and the number of constraints is $\mathcal{O}(p^{3})$.
+By Lenstra’s theorem [22] (and subsequent improvements by Kannan [21]), integer linear feasibility with a fixed number $d$ of variables can be solved in time $d^{\mathcal{O}(d)}\cdot\operatorname{poly}(L)$, where $L$ is the encoding length of the ILP.
+Here $d\leq p^{3}$ and $L=\operatorname{poly}(|S|+|C|+|D|+p)$, hence the feasibility test runs in $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|+|D|)$ time.
+If feasible, the resulting integer solution specifies a valid covering; otherwise, none exists.
 $\Box$
-To summarize, for a candidate covering size $m$ and ground set $\mathfrak{R}_{p}(S)$, we construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ and test, via exhaustive search (Lemma 4.12), whether $\mathfrak{R}_{p}(S)$ can be covered.
+To summarize, for a candidate covering size $m$ and ground set $\mathfrak{R}_{p}(S)$, we construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ and test, via a feasibility ILP (Lemma 4.12), whether $\mathfrak{R}_{p}(S)$ can be covered.
 If no covering is found, we conclude that $\mathfrak{R}_{p}(S)$ cannot be covered with $[m]$ (bad guess).
@@
 Theorem 4.13.
-There is a PTAS for the Double Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
+There is a $(1+\epsilon)$-approximation for the Double Prefix Covering problem that terminates in $\mathcal{O}(p^{\mathcal{O}(p^{3})}poly(|S|,\boldsymbol{OPT}))$, where $p=3/\epsilon^{2}$.
 Proof.
-Given an instance $S$ of Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+1/p)^{2}\leq 1+\epsilon$.
+Given an instance $S$ of Double Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+1/p)^{2}\leq 1+\epsilon$.
 Construct $\mathfrak{R}_{p}(S)$ by rounding up elements in $S$.
@@
   1. (i).
-    Construct $\mathfrak{R}_{p}(C)$ where $C=[m]$ as per Definition 4.3.
+    Construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ where $C=[m],D=[m-\beta]$ as per Definition 4.3.
   2. (ii).
-    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
+    Use the feasibility ILP (Lemma 4.12) to determine if $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(p^{O(p^{3})}poly(|S|,{\boldsymbol{br}^{*}})$.
 Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$.
@@
 Using Proposition 4.11, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)^{2}\boldsymbol{OPT}(S)$.
-The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$.
+The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(p^{\mathcal{O}(p^{3})}poly(|S|,{\boldsymbol{br}^{*}}))$ time, giving total complexity $p^{\mathcal{O}(p^{3})}poly(|S|,{\boldsymbol{br}^{*}}))$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$.
 $\Box$
-Combining Lemma 4.10 and Theorem 4.13 implies the following Corollary.
+EPTAS for Telephone Broadcasting in Melon graphs
+Given the equivalence between Telephone Broadcasting in melon graphs and the Double Prefix Covering problem, as established in Lemma 4.10 along with the EPTAS presented in Theorem 4.13 for the Double Prefix Covering problem, we can conclude the following.
 Corollary 4.14.
-There exists a PTAS for Telephone Broadcasting in $k$-path graphs.
-In particular, for a $k$-path graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
-5 Graphs of Bounded Bandwidth
-In this section, we present a polynomial algorithm that computes the optimal broadcasting time for graphs of bounded bandwidth.
-Intuitively, the bandwidth constraint limits the degree of $G$ (in contrast to other families such as $k$-cycle or $k$-path graphs), and a bandwidth ordering exposes a small “active window” of vertices at each step.
+There exists an EPTAS for Telephone Broadcasting in melon graphs.
+In particular, for a melon graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
+Multi-source Broadcasting in Melon Graphs
+The EPTAS of Corollary 4.14 extends naturally to the multi-source broadcasting setting.
+Suppose we are given an instance $I=(G,S)$ of multi-source broadcasting in $G$, where $S$ is the set of all sources, along with a fixed parameter $\epsilon>0$.
+Let $\boldsymbol{br}$ be a broadcast-time guess value.
+We describe a procedure that, when $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, outputs a valid broadcasting scheme that completes within $(1+\epsilon)\boldsymbol{br}+1$.
+Otherwise, it returns either Bad-Guess or a scheme with no guarantee on its broadcast time.
+As we do not know the optimal broadcast time, we iterate over all $\boldsymbol{br}\in[\lceil\log n\rceil,n]$ and take the scheme with the best broadcasting time among all returned schemes.
+This way, we can achieve broadcast time $(1+\epsilon){\boldsymbol{br}^{*}}(I)+1$.
+Preprocessing.
+For each non-endpoint source $s_{i}$ in the melon graph, we remove all vertices that lie within distance $\boldsymbol{br}$ of $s_{i}$ on their respective paths, excluding the $s$ and $t$.
+(The vertex $s_{i}$ is also removed.) If the endpoints are not initially source vertices, let $s_{ms}$ and $s_{mt}$ be the sources closest to $s$ and $t$, respectively, at distance $d_{\min}(s)$ and $d_{\min}(t)$.
+Let $G^{\prime}$ denote the remaining graph and $d_{\min}=\min(d_{\min}(s),d_{\min}(t))$.
+If $G^{\prime}$ is disconnected, or if $\max(d_{\min}(s),d_{\min}(t))>\boldsymbol{br}$, then there exists a vertex whose distance from every source exceeds $\boldsymbol{br}$, and we correctly return Bad-Guess.
+The preprocessing removes a window of at most $\boldsymbol{br}$ vertices around each non-center source.
+Note that all removed vertices can be informed within $\boldsymbol{br}+1$ rounds by broadcasting from their respective sources.
+Reduction to covering.
+Consider an instance $I^{\prime}=(G^{\prime},\{s,t\},\alpha)$ of Telephone Broadcasting, where $s$ receives the message at time 0 and $t$ at time $\alpha$.
+We associate $I^{\prime}$ with an instance $\mathcal{S}$ of the Double Prefix Covering problem, defined analogously to Lemma 4.10 for melon graphs.
+The only difference is that some elements of the ground multiset $\mathcal{S}$, which we refer to as partial path elements, correspond to partially-covered paths in $G^{\prime}$ rather than full paths.
+Each such path, depending on the endpoint that it is connected to, must be covered by exactly one of the endpoints (it cannot be covered by both covering sets).
+This modified covering instance can be solved in the same way as discussed in Section 4.2.
+In particular, the rounding steps remain unchanged.
+The only additional constraint arises in the final step (Lemma 4.12), where we ensure that each partial path element in the rounded ground set is covered by its corresponding rounded cover set.
+This produces a broadcast schedule for $I^{\prime}$ that completes within $(1+\epsilon)\,{\boldsymbol{br}^{*}}(I^{\prime})$ rounds.
+Combining the solutions. Overall, broadcasting in the original multi-source instance $G$ completes within
+$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(I^{\prime})\,\right\}.$
+The term $\boldsymbol{br}+1$ accounts for informing all vertices removed during the preprocessing step.
+The second term captures the time required to inform the closest endpoint and possibly $1$ extra round if the same source covers both endpoints (one of the endpoints can be informed through the other endpoint), which is $d_{\min}+1$ rounds, followed by the execution of the single-source broadcast within $I^{\prime}$.
+Suppose $br={\boldsymbol{br}^{*}}(I)$.
+Now, if $G^{\prime}$ has no vertices other than $s$ and $t$, then we have found a broadcasting schedule with $\boldsymbol{br}+1={\boldsymbol{br}^{*}}(I)+1$ rounds.
+Otherwise, in the optimal scheme, vertices in $G^{\prime}$ must be informed through the $s$ and $t$.
+This is because the vertices in $G^{\prime}$ are the ones that were not removed when forming $G^{\prime}$ and thus are at a distance further than $\boldsymbol{br}={\boldsymbol{br}^{*}}(I)$ from the sources in their respective cycles.
+Therefore, as the center cannot be informed earlier than round $d_{\min}$, we can write ${\boldsymbol{br}^{*}}(I)\geq d_{\min}+{\boldsymbol{br}^{*}}(I^{\prime})$.
+Hence, the broadcast time of our scheme is at most
+$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}\leq(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1.$
+5 Graphs of Bounded Cutwidth
+In this section, we give a polynomial-time algorithm for optimally solving the multi-source broadcasting problem on graphs of bounded cutwidth.
+Intuitively, the cutwidth constraint limits the degree of $G$ (in contrast to other families such as cycle-star or melon graphs), and a cutwidth ordering exposes a small “active window” of vertices at each step.
 This property enables a dynamic programming (DP) approach.
-For the remainder of this section, we assume that a bandwidth-$k$ ordering $(v_{1},\ldots,v_{n})$ of the vertices is given.
-For clarity of exposition, we also assume that the source of the broadcast is $v_{1}$; this restriction will be relaxed later.
+We assume throughout that a cutwidth-$k$ ordering $(v_{1},\ldots,v_{n})$ is given; all references to “before” and “after” are with respect to this ordering.
+Recall that such an ordering can be computed in fixed-parameter tractable time for constant $k$ [28].
 DP Table.
 We define a DP table with entries of the form $dp[i][s]$, where $i\in[0,n]$ and $s$ is a state.
-The value $dp[i][s]$ denotes the minimum broadcast time for informing the first $i$ vertices in the ordering while ending in configuration $s$.
-The state $s$ encodes all necessary information about the boundary vertices $B_{i}=\{v_{i-k+1},\ldots,v_{i}\}$, namely, how these vertices are informed and how they connect to the rest of a potential broadcast tree.
+The value $dp[i][s]$ denotes the minimum broadcast time for informing the first $i$ vertices in the cutwidth ordering while ending in state $s$.
+Define the set of boundary vertices as $B_{i}=\{v_{j}|j\leq i,v_{j}\text{ has at least a neighbor after }v_{i}\}$.
+Given that cutwidth is $k$, it holds that $|B_{i}|\leq k$.
+The state $s$ encodes all necessary information about the boundary vertices, namely, how these vertices are informed and how the vertices appearing after $v_{i}$ can connect to them to form a potential broadcast tree.
 Intuitively, a state describes the partial structure of a broadcast forest spanning boundary vertices together with their timing information.
-We fill the table row by row, starting from $i=k$ up to $i=n$. The last row corresponds to the entire graph; the minimum value among its states equals the optimal broadcast time.
+We fill out the table row by row, starting with $i=0$ and ending with $i=n$.
+The last row corresponds to the entire graph, and the minimum value among its states indicates the optimal broadcast time.
 Definition of states.
-Fix $i$, and let $u_{1},\ldots,u_{k}$ denote the boundary vertices.
-Let $Bf=\{x_{1},\ldots,x_{k}\}$ be the $k$ vertices immediately before $B_{i}$ and $Af=\{y_{1},\ldots,y_{k}\}$ be the $k$ vertices immediately after $B_{i}$.
-Together, the vertices $x_{1},\ldots,x_{k},\;u_{1},\ldots,u_{k},\;y_{1},\ldots,y_{k}$ form the active window.
-Every neighbor of a boundary vertex lies within this window.
+Fix $i$, and recall $B_{i}$ denotes the set of boundary vertices.
+We define the active window of $i$, denoted by $A_{i}$, as a set formed by the union of $B_{i}$ and all neighbors of vertices in $B_{i}$.
+Given a cutwidth-$k$ ordering, the degree of every vertex is at most $2k$.
+For an index $i$, the active window contains all vertices in $B_{i}$ (at most $k$ vertices) and their neighbors.
+Therefore, its total size is bounded by $k+2k^{2}$.
 A state $s$ encodes a potential partial broadcast tree consistent with the active window by storing the following information:
   • Parents.
-For each $u_{i}$, its parent $p(u_{i})$ (unless it is the source $v_{1}$).
-Each $p(u_{i})$ is chosen from at most $2k$ active vertices or a special “no parent” symbol, yielding $(2k+1)^{k}$ possibilities.
+For each $u\in B_{i}$, its parent $p(u)$ (unless it is a source vertex).
+Each $p(u)$ is chosen from at most $2k$ active vertices or a special “no parent” symbol for the sources, yielding $(2k+1)^{k}$ possibilities.
   • Ranks.
-For each $u_{i}$, its rank among the children of $p(u_{i})$, ordered by the rounds in which $p(u_{i})$ calls them.
-Since $p(u_{i})$ has at most $2k$ children, there are at most $2k$ options per vertex, and $(2k)^{k}$ total options.
+For each $u\in B_{i}$, its rank among the children of $p(u)$, ordered by the rounds in which $p(u)$ calls them.
+Since $p(u)$ has at most $2k$ children, there are at most $2k$ options per vertex, and $(2k)^{k}$ total options.
   • Outgoing calls.
-For each $u_{i}$, the active vertex that it calls in each round after being informed.
-For every round slot, there are $2k$ options (one of up to 2k12k-1 children or a null option if unused), giving at most $(2k)^{2k+1}$ possibilities per $u_{i}$.
+For each $u\in B_{i}$, the active vertex that it calls in each round after being informed.
+For every round slot, there are $2k+1$ options (one of up to $2k$ children or a null option if the slot is unused), giving at most $(2k+1)^{2k+1}$ possibilities per $u\in B_{i}$.
   • Inform times.
 The time at which the root of each subtree is informed.
-Since there are at most $k$ subtrees (each containing a boundary vertex), and each root may be at a round in $[n]$, this yields $n^{k}$ possibilities.
+Since there are at most $k$ subtrees (each containing a boundary vertex), and each root may be informed at a round in $[n]$, this yields $n^{k}$ possibilities.
 Proposition 5.1.
-If $S$ denotes the space of all states, then $|S|\;\leq\;(2k+1)^{k}\cdot(2k)^{k}\cdot(2k)^{k(2k+1)}\cdot n^{k}$, which is $\mathcal{O}(n^{k})$ for constant $k$.
+If $S$ denotes the space of all states, then
+$|S|$ $\;\leq\;f(k)\cdot n^{k},$
+$\text{\ where \ \ }f(k)$ $=(2k+1)^{k}\cdot(2k)^{k}\cdot(2k+1)^{k(2k+1)}.$
+In particular, $|S|=\mathcal{O}(n^{k})$ for constant $k$.
 For a state $s$, let $\ell(s)$ denote the latest time at which any active node in the broadcast subtrees of $s$ is informed.
 Note that the global broadcast time of a scheme realizing $s$ may exceed $\ell(s)$ if the last informed vertex lies outside the active window.
-Filling the DP table.
+Filling Out the DP table.
 A state $s$ is valid for step $i$ if its parent–child assignments among nodes in its broadcast forest respect the edges of $G$.
-Also, the outgoing calls in the state must match the rank and parent fields.
 For example, if $u_{1}$ and $u_{2}$ are not adjacent in $G$ but $u_{1}$ is the parent of $u_{2}$ in $s$, then $s$ is invalid at $i$.
+Furthermore, the inform times of all source vertices must be 0.
 For any invalid state at step $i$, we set $dp[i][s]=+\infty$.
-Base case. When $i=k$, any valid state represents a single broadcast tree rooted at $v_{1}$ (informed at time 0). For such states, we set $dp[k][s]=\ell(s)$.
+Base case. When $i=0$, the active window is an empty set. Therefore, there is only one null valid state $s_{\phi}$, for which $dp[0][s_{\phi}]=0$.
 Transitions.
-To compute $dp[i+1][s]$ for a valid state $s$, we look for states $s^{\prime}$ that are valid at step $i$ and compatible with $s$.
+To compute $dp[i+1][s]$ for a valid state $s$, we check states like $s^{\prime}$ that are valid at step $i$ and compatible with $s$.
 Compatibility means that information concerning the overlap between active states at steps $i$ and $i+1$ matches in both states: the inform times of shared vertices, their parent–child relations, and the directions and timings of edges incident to them must all agree.
-Further, indices shift in the natural way: for example, $u_{1}$ at step $i$ corresponds to $x_{k}$ at step $i+1$, and $u_{2}$ at step $i$ corresponds to $u_{1}$ at step $i+1$.
-Vertices $x_{1}$ and $y_{k}$ enter or leave the active window accordingly.
-See Figure 4 for an illustration.
 If $s^{*}$ is a compatible state with $s$ with minimum $dp[i][s^{*}]$, we set $dp[i+1][s]\;=\;\max\{dp[i][s^{*}],\ell(s)\}.$ We also record $\mathrm{previous}(s)=s^{*}$, which allows reconstruction of the optimal broadcast tree.
@@
 Intuitively, realizing state $s$ at $i+1$ requires extending from some compatible predecessor state at $i$, and the broadcast time is determined by whichever is larger: the time already incurred or the latest inform time among the active vertices.
-Example 9.
-In Figure 4, suppose that at step $i+1$, the best compatible predecessor of state $S_{2}$ is $S_{1}$ with $dp[i][S_{1}]=120$. Since $\ell(S_{2})=109$, we obtain $dp[i+1][S_{2}]=\max\{120,109\}$.
-[…]
-Figure 4: An illustration of states and their compatibility for a graph of bandwidth $k=3$.
-The bandwidth ordering is $(\ldots,x_{1},x_{2},x_{3},u_{1},u_{2},u_{3},y_{1},y_{2},y_{3},\ldots)$.
-The boundary vertices $u_{i}$ are highlighted in red.
-When computing DP values at step $i+1$, state $S_{2}$ is compatible with state $S_{1}$.
-Note that the shift in the active window implies, for example, that $u_{1}$ in $S_{1}$ at step $i$ becomes $x_{3}$ in $S_{2}$ at step $i+1$, $u_{2}$ becomes $u_{1}$, $x_{2}$ becomes $x_{1}$, and so on.
-Sources other than $v_{1}$.
-If the broadcast source is not $v_{1}$, then we consider all $O((2k)!)$ possible neighbor orderings in which the source $s$ informs its neighbors.
-For a fixed ordering, the problem splits into two independent subproblems: one for vertices preceding $s$ in the bandwidth ordering and one for vertices following it.
-The DP applies symmetrically to both sides.
-The only modification is in the base case: a state is valid only if it is consistent with the fixed ordering of $s$’s initial calls.
-Since $k$ is constant, this does not affect the asymptotic time complexity.
 Using an inductive argument, one can establish the correctness of the above approach.
 Theorem 5.2.
-Telephone Broadcasting can be solved optimally on graphs of bandwidth $k$ in time $f(k)\cdot n^{k^{2}+O(k)}$.
+Telephone Broadcasting can be solved optimally on graphs of cutwidth $k$ in time $f(k)^{2}\cdot n^{2k+1}$.
 In particular, for a constant $k$, there is a polynomial-time algorithm for computing an optimal broadcast schedule.
@@
 Setup.
-Recall that we have a fixed bandwidth-$k$ ordering $(v_{1},\dots,v_{n})$.
-As discussed, at step $i$ the DP keeps states $s$ that summarize the broadcast restricted to the active window $W_{i}:=\{x_{1},\dots,x_{k},u_{1},\dots,u_{k},y_{1},\dots,y_{k}\}$, where $B_{i}=\{u_{1},\dots,u_{k}\}$ is the set of boundary nodes.
+Recall that we have a fixed cutwidth-$k$ ordering $(v_{1},\dots,v_{n})$.
+As discussed, at step $i$, the DP table has states $s$ that summarize the broadcast restricted to the active window $A_{i}$ with boundary set $B_{i}$.
 Recall that, for a state $s$, $\ell(s)$ is the latest inform time among active vertices encoded by $s$.
-Correctness. We use induction to prove the following invariant:
-Inductive invariant.
-For every $i\geq k$ and every valid state $s$ at step $i$, the entry $dp[i][s]$ of the DP table is the minimum broadcast time of any feasible broadcast scheme for informing $v_{1},\ldots,v_{i}$ which ends up at state $s$.
-Base $i=k$.
-In the base case, the round in which each vertex of $v_{1},\ldots,v_{k}$ is informed is embedded into the state $s$.
-This is because they are all boundary vertices.
-Therefore, the minimum broadcast time over schedules realizing $s$ is exactly $\ell(s)$, so $dp[k][s]=\ell(s)$ as set by the algorithm.
-Step $i\to i+1$. Let $s$ be any valid state at step $i+1$. Choose a compatible predecessor $s^{*}$ at step $i$ that minimizes $dp[i][s^{*}]$.
+Correctness.
+Define the overlap signature of a state $s$ at step $i$ as the restriction $\pi_{i}(s)$ of $s$ to the overlap $A_{i}\cap A_{i+1}$: this includes (i) parent/child relations among vertices in the overlap, (ii) the directions/times of incident used edges, and (iii) their inform times.
+By definition, two states $s^{\prime}$ at $i$ and $s$ at i+1i{+}1 are compatible iff $\pi_{i}(s^{\prime})=\pi_{i}(s)$.
+We use induction to prove the following invariant:
+Claim 1 (Inductive invariant).
+For every $i\geq 0$ and every valid state $s$ at step $i$, the entry $dp[i][s]$ of the DP table is the minimum broadcast time of any feasible broadcast scheme for informing $v_{1},\ldots,v_{i}$ which ends up at state $s$ for the active set.
+Base $i=0$. As there are no vertices in this case, the answer for state $s_{\phi}$ is trivially 0 to inform an empty set.
+Step $i\to i+1$.
+Let $s$ be any valid state at step $i+1$.
+Choose a compatible predecessor $s^{*}$ at step $i$ that minimizes $dp[i][s^{*}]$.
 By the inductive hypothesis there is broadcast scheme $\mathcal{S}^{*}$ that realizes $s^{*}$ and informs $v_{1},\ldots,v_{i}$ in $dp[i][s^{*}]$ rounds.
-Extend $\mathcal{S}^{*}$ inside the active window to realize $s$ at step i+1i{+}1 (this is possible by compatibility and validity).
-The broadcast time of the extended scheme is $\max(dp[i][s^{*}],\ell(s))$, which is exactly the value assigned by the recurrence.
+Extend $\mathcal{S}^{*}$ inside the active window to realize $s$ at step i+1i{+}1 (this is possible by compatibility of $s$ and $s^{*}$ and their validity).
+The broadcast time of the extended scheme is $\max\{dp[i][s^{*}],\ell(s)\}$, which is exactly the value assigned by the recurrence.
 Hence, dp[i+1][s]dp[i{+}1][s] upper-bounds the true optimum among schedules realizing $s$.
@@
 By the inductive hypothesis, $dp[i][s^{\prime}]$ is no more than the broadcast time of $\mathcal{S}$ for informing $v_{1},\ldots,v_{i}$.
-Since realizing $s$ from $s^{\prime}$ cannot finish later than $\max\{dp[i][s^{\prime}],\ell(s)\}$, we get $dp[i{+}1][s]\leq\max\{dp[i][s^{\prime}],\ell(s)\}\leq\tau.$ Thus, dp[i+1][s]dp[i{+}1][s] lower-bounds the optimum as well.
-This proves the inductive invariant.
+Since realizing $s$ from $s^{\prime}$ cannot finish later than $\max\{dp[i][s^{\prime}],\ell(s)\}$, we get $dp[i{+}1][s]\leq\max\{dp[i][s^{\prime}],\ell(s)\}\leq\tau.$ Thus dp[i+1][s]dp[i{+}1][s] lower-bounds the optimum as well.
+This proves Claim 1.
 We conclude that taking the minimum $dp[n][s]$ over valid states $s$ yields the optimal broadcast time on $G$.
-Sources other than $v_{1}$.
-If the source is $s\neq v_{1}$, we enumerate the $O((2k)!)$ possible orders of $s$’s calls (only neighbors within distance $k$ in the bandwidth ordering matter).
-For each fixed order, the same inductive argument applies independently to the two sides of $s$, and we take the best over the constant many orders.
-This preserves optimality.
 Running time.
 Let $S$ be the set of all states.
-By Proposition 5.1, we have $|S|\leq(2k{+}1)^{k}\cdot(2k)^{k}\cdot(2k)^{k(2k+1)}\cdot n^{k}\;=\;g(k)\,n^{k}.$ That is, the size of the DP table would be at most $g(k)n^{k+1}$.
-For every $i$ and $s$, comparing every state $s^{\prime}$ of the previous iteration (i.e.
-i1i-1) to find out whether they are compatible would take $O(|S|.n)$ time.
-Therefore, the cost of computing dp for all states is $(g(k))^{2}\,n^{2k+2}$.
-Overall, this yields a total time $f(k)\cdot n^{O(k)}$, for some function $f(k)$ independent of $n$.
-Note that the extra $O((2k)!)$ factor for non-$v_{1}$ sources is absorbed into $f(k)$ since $k$ is constant.
-This matches the claimed bound.
+By Proposition 5.1, we have $|S|\leq(2k{+}1)^{k}\cdot(2k)^{k}\cdot(2k+1)^{k(2k+1)}\cdot n^{k}\;=\;f(k)\,n^{k}.$ That is, the size of the DP table would be $|S|n=f(k)n^{k+1}$.
+To set each index of the DP table, we iterate over all possible states, which takes $\mathcal{O}(|S|)$ time.
+Overall, this yields a total time of $\mathcal{O}(f(k)^{2}\,n^{2k+1})$.
 $\Box$
+Note that the above dynamic programming approach is formulated for instances with multiple sources, and therefore Theorem 5.2 extends to the multi-source setting.
+6 Improved Approximation Algorithm for Split Graphs
+In this section, we present a polynomial-time algorithm that finds a broadcasting scheme which completes within $1.76{\boldsymbol{br}^{*}}$ rounds for split graphs.
+Consider an instance $(G,s)$ of Telephone Broadcasting in an input split graph $G=(V,E)$ whose vertex set can be partitioned into a clique $C$ and an independent set $I$.
+Throughout, we use $B$ to denote the bipartite graph induced by the edges incident to $I$.
+We begin by introducing the necessary preliminaries and then describe our algorithm.
+Minimum-degree cover.
+An important concept related to broadcasting on split graphs is the minimum-degree cover, which also plays a key role in our algorithm.
+A cover of $B$ is a subset $M$ of edges of $B$ such that every vertex $u\in I$ is incident to exactly one edge of $M$.
+For a clique vertex $v\in C$, let $\deg_{M}(v)$ denote the degree of $v$ under $M$, and define the degree of $M$ as $\Delta(M):=\max_{v\in C}\deg_{M}(v)$.
+Let $d^{*}(B):=\min\{\Delta(M)\mid M\text{ is a cover of }B\}$.
+We refer to a cover attaining degree $d^{*}(B)$ as a minimum-degree cover of $B$.
+When the bipartite graph $B$ is clear from the context, we write $d^{*}$ instead of $d^{*}(B)$ (See Figure 4).
+The value $d^{*}$ can be computed in polynomial time using a straightforward max-flow formulation.
+Introduce a source $s^{\prime}$, a sink $t^{\prime}$, connect $s^{\prime}$ to each $u\in C$ with capacity $D$, connect each $(u,v)\in E(B)$ with capacity $1$, and connect each $v\in I$ to $t^{\prime}$ with capacity $1$.
+The smallest value $D$ for which a flow of value $|I|$ exists is exactly $d^{*}$.
+Minimum-overload covers.
+Fix parameters $\tau_{\ell}<\tau_{u}$.
+For a cover $M$, define the overload of a clique vertex $v\in C$ by $\mathrm{ov}_{M}(v):=\max\{0,\deg_{M}(v)-\tau_{\ell}\},$ and define the total overload as $\mathrm{OV}(M):=\sum_{v\in C}\mathrm{ov}_{M}(v).$ Among all covers $M$ with $\Delta(M)\leq\tau_{v}$, let $M^{\mathrm{ov}}$ be one minimizing $\mathrm{OV}(M)$, and let $R:=\mathrm{OV}(M^{\mathrm{ov}})$.
+See Figure 4 for an illustration.
+Among all covers of $B$, the one with the minimum-overload cover can be found via a reduction to the min-cost flow, which can be solved optimally in polynomial time (see, e.g.,  [1]).
+For the reduction, we construct a network flow from $B$ as follows: $s^{\prime}$ is connected to every vertex $u\in C$ with two parallel edges: one of capacity $\tau_{\ell}$ and cost 0, and one of capacity $\tau_{u}-\tau_{\ell}$ and cost $1$; each edge $(u,v)\in E(B)$ has capacity $1$ and cost 0; each vertex $v\in I$ is connected to the sink with an edge of capacity $1$ and cost 0.
+Sending one unit of flow through the costly edge corresponds to one unit of overload.
+Minimizing the total cost yields a cover minimizing overload.
+$u_{1}$$u_{2}$$u_{3}$
+(a) Cover $M$
+$u_{1}$$u_{2}$$u_{3}$
+(b) Cover $M^{\prime}$
+Figure 4: Two minimum-degree coverings of $B$ with $d^{*}=3$.
+When $\tau_{\ell}=2$, we have $OV(M)=2$ (both $u_{1}$ and $u_{2}$ have overload 1), while $OV(M^{\prime})=1$ (only $u_{1}$ has overlaod 1).
+The Basic algorithm. We begin with a warm-up approximation algorithm of [16], which we call Basic. The algorithm works as follows:
+  1. (i).
+    Broadcast within the clique $C$ until all clique vertices are informed.
+  2. (ii).
+    Use a minimum-degree cover $M$ of $B$ to inform the vertices in $I$: each $v\in C$ informs its assigned neighbors one per round.
+Step (1) completes in at most $\lceil\log|C|\rceil$ rounds.
+Step (2) requires $\Delta(M)=d^{*}$ additional rounds.
+Thus, $T_{\textsc{Simple}}\leq\lceil\log|C|\rceil+d^{*}.$ Both $\lceil\log|C|\rceil$ and $d^{*}$ are lower bounds on $\boldsymbol{br}^{*}$}: the former is necessary to inform all clique vertices, while the latter follows since each vertex of $I$ must be informed by some neighbor in $C$, and each clique vertex can inform at most one new vertex per round.
+Hence, Basic is a $2$-approximation.
+FastSplitBroadcast: Improved Approximation for Split Graphs
+Our approximation algorithm, named FastSplitBroadcast, refines Simple by distinguishing clique vertices according to their load in the minimum-overload cover.
+Intuitively, the vertices with high overload correspond to unavoidable congestion in any broadcast schedule and are handled in a separate phase to be informed earlier, thereby improving the approximation ratio.
+For that, we compute a cover that keeps the maximum load below a prescribed threshold while minimizing the extent to which the assignment exceeds a smaller “soft cap.” In particular, FastSplitBroadcast proceeds as follows (see Algorithm 1):
+0: Split graph $G=(C\cup I,E)$ with source $s$
+1: Compute a minimum-degree cover of $B$ and let $d^{*}$ be its degree.
+2: Set thresholds $\tau_{\ell}=\alpha d^{*}$, $\tau_{m}=(\alpha+\epsilon)d^{*}$, and $\tau_{u}=\beta d^{*}$ (for some $\alpha\leq 1\leq\beta,$ and small $\epsilon>0$).
+3: Compute a minimum-overload cover $M^{\mathrm{ov}}$ with cap $\tau_{u}$ and overload parameter $\tau_{\ell}$.
+4: Let $H:=\{u\in C:\deg_{M^{\mathrm{ov}}}(u)\geq\tau_{m}\}$ (high-load vertices) and
+$L:=C\setminus H$ (low-load vertices).
+5: Form the following broadcast scheme Class:
+  • Ensure that at least one vertex of $H$ and one vertex of $L$ are informed (within two rounds).
+  • Broadcast within $H$ until all vertices of $H$ are informed.
+  • Broadcast from vertices of $H$ to their neighbors in $I$ assigned by $M^{\mathrm{ov}}$.
+  • In parallel, broadcast within $L$ and then from $L$ to their neighbors in $I$ assigned by $M^{\mathrm{ov}}$.
+6: Between Basic and Class, return the faster scheme.
+Algorithm 1 SplitBroadcast (single source)
+Step 1: Compute a minimum-degree cover of $B$ and let $d^{*}$ be its degree.
+Step 2: Set thresholds $\tau_{\ell}=\alpha d^{*}$, $\tau_{m}=(\alpha+\epsilon)d^{*}$, and $\tau_{u}=\beta d^{*}$ for some $\alpha\leq 1\leq\beta,$ that will be decided later, and a small $\epsilon>0$.
+Step 3: Compute a minimum-overload cover.
+Use the min-cost flow reduction described earlier to compute the minimum overload cover of $B$ with parameters $\tau_{\ell}$ and $\tau_{u}$ in polynomial time.
+Step 4: Classify clique vertices.
+Using $M^{\mathrm{ov}}$, classify clique vertices as high-load if they are assigned at least $\tau_{m}=(\alpha+\epsilon)d^{*}$ neighbors, and low-load otherwise.
+Let $H$ be the set of high-load vertices and $L=C\setminus H$.
+Step 5: Form the broadcast scheme Class as follows
+  • Step 5(a): Seeding.
+Within two rounds, ensure that at least one vertex in $H$ and at least one vertex in $L$ becomes informed.
+It takes one round to inform one vertex $u$ in $C$ (if $s\in I$), and another round to inform another vertex in $C$ which belongs to a different group than $u$.
+  • Step 5(b): Two parallel broadcast processes.
+From this point on, run two broadcast processes in parallel.
+First, broadcast within the clique subgraph induced by $H$ until all vertices in $H$ are informed, and then each $u\in H$ informs its assigned independent neighbors (those $v\in I$ with $(u,v)\in M^{\mathrm{ov}}$) one per round.
+This part takes at most $\lceil\log|H|\rceil+\tau_{u}$ rounds after the seeding step.
+    In parallel, we broadcast within the clique subgraph induced by $L$ until all vertices in $L$ are informed, and then each $u\in L$ informs its assigned independent neighbors one per round.
+Since every low-load vertex has at most $\tau_{m}$ assigned neighbors, this part takes at most $\lceil\log|C|\rceil+\tau_{m}$ rounds after the seeding step.
+Step 6: Among Basic and Class, return the scheme that has a smaller broadcast time.
+Analysis of FastSplitBroadcast.
+We analyze the broadcast time of the algorithm.
+In what follows, we let $n$ denote $|C|$; recall that $d^{*}$ denotes the degree of a minimum-degree cover of $B$, and $M^{\mathrm{ov}}$ is a minimum-overload cover with parameters $\tau_{\ell}=\alpha d^{*}$, $\tau_{m}=(\alpha+\epsilon)d^{*}$, and $\tau_{u}=\beta d^{*}$.
+Let $H$ denote the set of clique vertices whose load in $M^{\mathrm{ov}}$ is at least $\tau_{m}$, and let $n_{h}=|H|$.
+Upper bound on the algorithm.
+After two seeding rounds, the algorithm Class runs two broadcast processes in parallel within $H$ and $L$.
+Broadcasting within $H$ takes at most $\lceil\log n_{h}\rceil$ rounds, followed by at most $\tau_{u}$ rounds to inform the independent vertices assigned to $H$ by $M^{\mathrm{ov}}$.
+Similarly, broadcasting within $L=C\setminus H$ takes at most $\lceil\log n\rceil$ rounds, followed by at most $\tau_{m}$ rounds to inform their assigned neighbors.
+Hence, the broadcast time of the algorithms satisfy
+$\boldsymbol{br}({\mathrm{CLASS}})\;$  $\leq\;2+\max\!\left\{\lceil\log n_{h}\rceil+\beta d^{*},\;\lceil\log n\rceil+(\alpha+\epsilon)d^{*}\right\},$ (6)
+$\boldsymbol{br}({\text{{FastSplitBroadcast}}})$ $=\min\{\boldsymbol{br}({\text{{Basic}}}),{\boldsymbol{br}(\text{{Class}}})\}.$ (7)
+Lower bounds for $\boldsymbol{br}^{*}$}.
+Throughout the analysis, we make use of the following lemma, which states that there exists an optimal broadcast tree in which every vertex in II—except possibly the source, if it belongs to II—appears as a leaf.
+uuvvzzAABBCCxxyy
+(a) Broadcast tree before the transformation
+uuvvzzAABBCCxxyy
+(b) Broadcast tree after the transformation
+Figure 5: An example showing how a broadcast scheme with a vertex $v\in I$ as a non-leaf can be transformed, without increasing the broadcast time, to eliminate non-leaf vertices in the independent set.
+Lemma 6.1.
+Let $G=(C\cup I,E)$ be a split graph with clique $C$ and independent set $I$, and let $s\in V(G)$ be the source vertex.
+There exists an optimal broadcasting scheme in which every non-source vertex of $I$ is a leaf in the broadcasting tree.
+Proof.
+Let $T$ be an optimal broadcast tree.
+Suppose $T$ contains a vertex $v\in I$ that is a non-source internal node, and let $z$ be the first vertex that $v$ informs.
+We show how to modify $T$, without increasing its broadcast time, so that $v$ has one fewer child.
+By repeating this transformation, we eventually obtain a broadcast tree in which every vertex in $I$ is either a leaf or the root, as desired.
+Modifying $T$: remove the edge $v\to z$ and instead have $u$ call $z$ at the time it originally called $v$, then call $v$ one step later.
+That is, $z$ takes $v$’s former slot in $u$’s calling sequence.
+Moreover, $z$ informs $v$ as its first child.
+See Figure 5 for an illustration.
+We verify that the broadcast time does not increase
+  • $z$ and its subtree: $z$ is informed one step earlier, so every vertex in its subtree is informed earlier.
+  • $v$ and its remaining children: $v$ is informed one step later, but has lost its first child $z$.
+Each remaining child of $v$ (originally called at times $t_{v}+2,t_{v}+3,\ldots$) is now called at $(t_{v}+1)+1,(t_{v}+1)+2,\ldots$ — the same times as before.
+  • Other vertices receive at the same time as before.
+Therefore, the modified tree remains optimal. $\Box$
+In light of the above lemma, we will consider an optimal cover, defined as a covering of $B$ induced by an optimal broadcast tree in which all vertices of $I$ appear only as leaves (or possibly the root) and are covered by their neighbors.
+This assumption is used in the proof of the following theorem, which is the main result of this section.
+Theorem 6.2.
+FastSplitBroadcast gives an approximation factor of 1.76 for Telephone Broadcasting in split graphs.
+Proof. As discussed before, we have ${\boldsymbol{br}^{*}}\geq\max\{\lceil\log n\rceil,d^{*}\}$. We use a case analysis to complete the proof.
+Case 1: Suppose the optimal cover has degree at least $\beta d^{*}$. This implies that $d^{*}\leq{\boldsymbol{br}^{*}}/\beta$. In this case, Simple completes in
+$\boldsymbol{br}(\textsc{Basic})\leq\lceil\log n\rceil+d^{*}\;\leq\;(1+1/\beta){\boldsymbol{br}^{*}}.$ (8)
+Since FastSplitBroadcast returns the better scheme between Basic and Class, its approximation factor in this case is at most $(1+1/\beta)$.
+Case 2: Conisder otherwise.
+Therefore, $d^{*}>{\boldsymbol{br}^{*}}/\beta$.
+Recall that, in the broadcast scheme of Class, the last leaf with a parent in $H$ is informed no later than $\lceil\log n_{h}\rceil+\beta d^{*}$.
+In this case,
+$2+\lceil\log n\rceil+(\alpha+\epsilon)d^{*}\;\leq\;(1+\alpha+\epsilon)\,{\boldsymbol{br}^{*}}+2,$ (9)
+By optimality of $M^{\mathrm{ov}}$, the total overload $R$ of $M^{\mathrm{ov}}$ with respect to $\tau_{\ell}$ and $\tau_{u}$ is at most the overload of the optimal cover.
+Moreover, in the optimal cover, each clique vertex whose load is at least $\tau_{\ell}$ contributes at most $\tau_{u}-\tau_{\ell}=(\beta-\alpha)d^{*}$ units of overload.
+It follows that in the optimal cover, there are at least $\frac{R}{(\beta-\alpha)d^{*}}$ clique vertices of load at least $\tau_{\ell}$.
+Informing these vertices and their neighbors in $I$ defines a lower bound for the optimal broadcast time.
+That is,
+${\boldsymbol{br}^{*}}\geq\lceil\log\frac{R}{(\beta-\alpha)d^{*}}\rceil+\alpha d^{*}\geq\log\frac{R}{d^{*}}+\alpha d^{*}-\log(\beta-\alpha).$
+On the other hand, every vertex in $H$ contributes at least $\tau_{m}-\tau_{\ell}=\epsilon d^{*}$ units of overload in the broadcast scheme of Class, implying $n_{h}\leq\frac{R}{\epsilon d^{*}}$.
+Substituting this bound on $n_{h}$ into the upper bound on $T_{\mathrm{ALG}}$ yields
+$2+\lceil\log n_{h}\rceil+\beta d^{*}$ $\leq\log\frac{R}{\epsilon d^{*}}+\beta d^{*}+3$
+$=\log\frac{R}{d^{*}}+\beta d^{*}+3-\log\epsilon$
+$\leq\;\frac{\beta}{\alpha}\,{\boldsymbol{br}^{*}}+3-\log\epsilon-\frac{\beta}{\alpha}\log(\beta-\alpha)$
+$=\frac{\beta}{\alpha}\,{\boldsymbol{br}^{*}}+\mathcal{O}(1).$ (10)
+Therefore, for the broadcast time of Class, we can conclude
+$\boldsymbol{br}(\textsc{Class})$ $=2+\max\!\left\{\lceil\log n_{h}\rceil+\beta d^{*},\;\lceil\log n\rceil+(\alpha+\epsilon)d^{*}\right\}$ (by Equation 7)
+$\leq\max\{(1+\alpha+\epsilon){\boldsymbol{br}^{*}},\frac{\beta}{\alpha}{\boldsymbol{br}^{*}}+\mathcal{O}(1)\}$ (by Equations 9 and 10)
+To conclude, the broadcast time of FastSplitBroadcast is at most $(1+1/\beta){\boldsymbol{br}^{*}}$ in Case 1 (Equation 9) and at most $\max\{(1+\alpha+\epsilon){\boldsymbol{br}^{*}},\frac{\beta}{\alpha}{\boldsymbol{br}^{*}}+\mathcal{O}(1)\}$ in Case 2.
+This gives an approximation factor of $\max\{1+1/\beta,1+\alpha+\epsilon,\beta/\alpha\}$ for FastSplitBroadcast.
+For $\alpha=0.754,\beta=1/\alpha,$ and small $\epsilon$, this would be come $1/\alpha^{2}\approx 1.76$.
+$\Box$
+Extension to multi-source telephone broadcasting
+In the multi-source setting, we first ensure that a set $U$ of clique vertices—consisting of clique sources and clique vertices matched to sources in the independent set—is informed in the first round.
+We then restrict attention to the remaining bipartite graph and apply the same covering-based strategy as in the single-source case.
+The rest of the algorithm and its analysis carry over verbatim, with the number of initially informed clique vertices $|U|$ effectively reducing the broadcast depth by $\lfloor\log|U|\rfloor$.
+We now describe the extension in more detail.
+Let $I_{s}$ and $C_{s}$ denote the sets of source vertices in $I$ and $C$, respectively.
+If $I_{s}$ is nonempty, we first compute a maximal matching between $I_{s}$ and vertices in $C\setminus C_{s}$.
+Let $U\subseteq C$ be the set of clique vertices that are either sources themselves or matched to a source in $I_{s}$.
+Consequently, all vertices in $U$ are informed by the end of round $1$.
+Let $B^{\prime}$ denote the bipartite graph induced by the vertex sets $C$ and $I\setminus I_{s}$.
+After round 1, the vertices in $U$ have the message and must inform the remaining uninformed vertices in $C\cup I$.
+From this point onward, the algorithm proceeds analogously to the single-source setting.
+In particular, we compute the minimum-degree cover and the minimum-overload cover with respect to $B^{\prime}$.
+The parameters $\alpha$ and $\beta$ are defined in the same way as before, and FastSplitBroadcast selects the better of the two strategies Basic and Class.
+The procedures Basic and Class are also defined analogously.
+The only modification is that, in the second round, Class ensures that at least $|U|$ vertices from each of the sets $L$ and $H$ are informed.
+Subsequently, vertices in $L$ and $H$ are informed in parallel, requiring at most $\lceil\log(n/|U|)\rceil$ and $\lceil\log(n_{H}/|U|)\rceil$ additional rounds, respectively.
+The analysis is identical to the single-source case, with the only difference that $n$ is replaced by $n/|U|$ in the cost analysis of both FastSplitBroadcast and the optimal broadcast scheme.
 7 Concluding Remarks
 In this work, we advanced the study of Telephone Broadcasting by establishing new hardness results and developing approximation schemes for specific graph families.
-On the hardness side, we resolved long-standing questions by proving NP-hardness for $k$-cycle and $k$-path graphs.
-On the algorithmic side, we presented polynomial-time approximation schemes (PTAS) for Telephone Broadcasting in both $k$-cycle and $k$-path graphs.
-We also introduced an exact algorithm for graphs of bounded bandwidth, thereby extending the tractable frontier of the problem.
-As future work, it would be interesting to adapt the integer covering formulations to other broadcasting models or broader sparse graph classes.
-In particular, for series–parallel graphs, our hardness result via $k$-path graphs rules out tractability, yet, to the best of our knowledge, no constant-factor approximation is currently known.
+On the hardness side, we proved NP-hardness for very restricted families of cycle-star and melon graphs.
+On the algorithmic side, we presented efficient polynomial-time approximation schemes (EPTASs) for both cycle-star and melon graphs; to the best of our knowledge, these are the first PTAS-type results for Telephone Broadcasting.
+We also introduced an optimal, polynomial-time algorithm for graphs of bounded cutwidth, thereby extending the tractable frontier of the problem.
+Finally, we developed a polynomial-time approximation algorithm for split graphs that improves upon the previously known factor–2.
+A possible direction for future work is to improve the approximation factor for split graphs.
+In particular, it remains an open question whether Polynomial Time Approximation Schemes (PTASs) exist for these graph classes.
+The major open problem in this domain is determining whether a constant-factor approximation exists for general graphs.
+While progress on this question has been slow, the algorithmic ideas developed in this work may be applicable to broadcasting in other families of sparse graphs.
+For example, it would be interesting to adapt the integer covering formulations to other broadcasting models or broader sparse graph classes.
+For series–parallel graphs, our hardness result via melon graphs rules out tractability, yet, to the best of our knowledge, no constant-factor approximation is currently known.